### v3

# Memorization Over Reasoning? Exposing and Mitigating Verbatim Memorization in Large Language Models’ Character Understanding Evaluation

Abstract
Recently, Large Language Models (LLMs) have shown impressive performance in character understanding tasks, such as analyzing the roles, personalities, and relationships of fictional characters. However, the extensive pre-training corpora used by LLMs raise concerns that they may rely on memorizing popular fictional works rather than genuinely understanding and reasoning about them. In this work, we argue that “gist memory”—capturing essential meaning—should be the primary mechanism for character understanding tasks, as opposed to “verbatim memory”—exact match of a string. We introduce a simple yet effective method to mitigate mechanized memorization in character understanding evaluations while preserving the essential implicit cues needed for comprehension and reasoning. Our approach reduces memorization-driven performance on popular fictional works from 96% accuracy to 72% and results in up to an 18% drop in accuracy across various character understanding tasks. These findings underscore the issue of data contamination in existing benchmarks, which often measure memorization rather than true character understanding.
Memorization Over Reasoning? Exposing and Mitigating Verbatim Memorization in Large Language Models’ Character Understanding Evaluation
Yuxuan Jiang , Francis Ferraro Department of Computer Science and Electrical Engineering University of Maryland, Baltimore County yuxuanj1@umbc.edu
1 Introduction
Character understanding requires the models to grasp the characters’ personalities and traits while reading a story, which aims to evaluate the models’ capability to comprehend personalities, goals, roles, and relationships of various fictional characters Yu et al. (2023a); Vallurupalli et al. (2024); Shen et al. (2024); Brahman et al. (2021). However, the impressive performance of LLMs on such benchmarks Li et al. (2023); Stammbach et al. (2022); Goyal et al. (2023) raises the question: does it reflect genuine understanding or merely strong memorization Chang et al. (2023a)? For example, in the TV show speaker guessing task shown in Figure  1, the model is expected to identify Sheldon based on its memory of the character from previous episodes (e.g., preparing lessons for Howard). However, if the model simply memorizes the exact words, it could make predictions based on a very short sentence by directly extracting the speaker information from it.
To truly stimulate a model’s character understanding ability, it is crucial to understand how the model may be generating its responses. This can be especially challenging given the extensive and opaque pre-training corpora used by LLMs, making it particularly severe for character understanding tasks, as many benchmarks are derived from well-known and widely consumed fiction or scripts Sang et al. (2022a). We hypothesize that this can lead to a form of memorization called “verbatim memory,” which encodes precise details and remembers every word exactly Huang et al. (2024); Ippolito et al. (2022)—rather than reasoning over recalled overarching patterns (what is referred to as “gist memory” Reyna and Brainerd (1998)).
Figure 1: This figure illustrates the distinct mechanisms of gist and verbatim memory processes in the speaker-guessing task for a script. Verbatim memory (bottom right) aims to recall as much specific information as possible, including who said each sentence. If the memorization is not precise enough, the model fails to provide the correct answer. In contrast, gist memory (top right) focuses on extracting core information, such as character relationships, personalities, and key events. By relying on gist memory, the model is encouraged to engage in reasoning rather than purely relying on memorization.
Motivated by these concepts, we propose an approach that alters elements such as character names, place names, and specific notations in the sentences while preserving the relationships between characters, significant events, and other information we want the model to utilize. We hypothesize that these modifications can reduce the model’s tendency to take shortcuts by relying on verbatim memory, and instead encourages the model to rely on gist memory to stimulate its reasoning ability for character understanding tasks. We tested our method on six key types of character understanding tasks and obtained an interesting finding: When the difficulty of using verbatim memory increases, the model is more likely to resort to reasoning to compensate.
Our work provides a simple approach to ensure a fairer benchmark for character understanding tasks on large models that already have strong memorization abilities of fictional works. It offers an alternative method compared to more costly solutions, such as developing data from obscure or very new books that the model has not encountered Ngo et al. (2024), or using techniques like machine unlearning to forget original training texts Shaik et al. (2024).
In summary, we provide an outline of the paper and summarize our key contributions:
  • To determine whether the memory types used in character understanding tasks align with our expectations, we draw on human cognitive science to define gist and verbatim memory in the context of character understanding tasks (s 2).
  • We benchmarked the mechanism memorization ability of over 100 sources of fictional works (s 4.1).
  • We measured the verbatim memory of large models for hundreds of different source fiction works and tested how different keywords affect the verbatim memory of these models (s 4).
  • We demonstrated that even for the same factor, such as a person’s name, different disturbances, such as gender or cultural interference, can lead to significantly different verbatim memory effects in large models (s 5).
  • We compared the different effects of the model using verbatim and gist memory across six different test sets and found that the model’s accuracy decreased by 0.7% to 17.2% (s 6).
2 Related Task
Among Large Language Models (LLMs)’ various capabilities Li et al. (2024b, d); Zhang et al. (2024); Li et al. (2024c), the reasoning abilities Tong et al. (2023); Yang et al. (2023); Xiong et al. (2024b); Tong et al. (2024); Xiong et al. (2024a) of LLMs Li et al. (2024a); Tan et al. (2024) have garnered significant attention, but their memory has also been shown to play a crucial role in influencing benchmark reasoning performance Zhou et al. (2023). This influence is complex: data contamination can inflate performance metrics Khandelwal et al. (2020); Blevins and Zettlemoyer (2022); Carlini et al. (2023), while knowledge conflict can reduce them Su et al. (2024); Wang et al. (2024); Xie et al. (2024b); Xu et al. (2024). Recent efforts have aimed to mitigate the impact of memory on reasoning benchmarks to establish more reliable evaluations, particularly for reasoning tasks like numerical reasoning, commonsense reasoning, or logical reasoning, where memory requirements are less significant and easier to define Xie et al. (2024a); Sainz et al. (2023); Li et al. (2024e). However, in character understanding tasks, the diversity of task formats and scenarios means that memory requirements vary significantly, making it challenging to establish a universal standard to distinguish between acceptable and unacceptable use of memory. We propose leveraging gist memory to reduce LLMs’ reliance on precise memorization. This approach encourages models to utilize reasoning capabilities more effectively in solving character understanding tasks, aligning them more closely with the original intent behind these benchmarks.
Character Understanding
Character understanding has made significant contributions to advancing story comprehension in psychological, literary, and educational research  Bower and Morrow (1990); McKee (2010); Paris and Paris (2003). This multifaceted field encompasses a variety of tasks, including character recognition, coreference resolution, summarization, role detection, goal understanding, personality analysis, and question answering Sang et al. (2022b); Bamman et al. (2020); Chen et al. (2022); Frermann et al. (2018); Vallurupalli et al. (2024); Yu et al. (2023b). Recent studies have shown promising performance by large language models (LLMs) on character understanding tasks Li et al. (2023); Stammbach et al. (2022); Goyal et al. (2023). However, these studies often overlook the potential influence of memorization on model performance. Fictional works, including movie and TV show scripts, famous novels, and news texts, serve as primary sources for narrative understanding Sang et al. (2022a); Tapaswi et al. (2016); Caselli and Vossen (2017). These texts, much like copyrighted books Chang et al. (2023a), are prone to strong memorization. In our work, we measured the verbatim capabilities of LLMs on these sources, as well as the potential inflation of benchmark results caused by overly strong verbatim memory. Our approach also mitigates this effect.
Verbatim and Gist Memory
Cognitive science defines two types of memory: verbatim memory and gist memory Brainerd and Reyna (2002); Reyna and Brainerd (1998). Verbatim memory recalls the exact wording of a sentence, such as "The black-and-white cat sat on a soft, red mat in the living room," retaining every detail about the cat’s color, the mat’s texture and color, and the location. In contrast, gist memory captures the essential meaning, like "A cat was sitting indoors," focusing on the core idea—the presence of a cat and its general location—while discarding details such as the cat’s color or the specific description of the mat. This highlights how verbatim memory retains all specific details, while gist memory abstracts the key information by ignoring less critical details. Cognitive science research suggests that humans primarily rely on gist memory for reasoning, and these two types of memory are independent, rather than hierarchical Brainerd and Reyna (1992). In recent studies on LLM memorization Carlini et al. (2023), the extraction of strings from a larger text is defined as a form of verbatim memory. Their solution primarily focuses on generating entirely new test samples that the model has never encountered Xie et al. (2024a). However, for character understanding tasks, using obscure or newly released datasets that the model hasn’t seen comes with high costs, and the quality and acquisition costs of the texts are difficult to control. Our work demonstrates that using gist memory provides a simple and effective way to alleviate this issue.
3 Data and Experiment Setup
In our experiments, we utilized two distinct datasets: (1) Benchmark Datasets: six pre-existing datasets designed to evaluate various character understanding tasks, and (2) Fictional Works Dataset: a newly constructed dataset consisting of script excerpts sampled directly from the original texts used to create the benchmark datasets. This dataset, drawn from 40 diverse sources—including TV shows, novels, and movies—was specifically designed to investigate GPT’s familiarity with the original texts underlying the benchmark datasets.
3.1 Benchmark Datasets
We evaluated the interplay between memory and reasoning across six types of character understanding tasks. The datasets used are detailed in this section, while task descriptions are provided in Section 6.1.
Datasets
For coreference resolution , we use the latest released version of the MovieCoref dataset  [^1] https://github.com/usc-sail/mica-character-coref. For character guessing, we adopt the TVSHOWGUESS dataset  [^2] https://github.com/YisiSang/TVSHOWGUESS to conduct experiments. For personality understanding, we utilize the PERSONET dataset  [^3] https://github.com/Gorov/personet_acl23. For role detection, we use the CSI dataset  [^4] https://github.com/EdinburghNLP/csi-corpus. For open-domain question answering, we conduct experiments using the FriendsQA dataset  [^5] https://github.com/emorynlp/FriendsQA. For summarization task, we adopt the SummScreen dataset  [^6] https://github.com/mingdachen/SummScreen. We follow all the data preprocessing, development, and testing separation provided by the original datasets. Tasks and datasets details are in appendix G.
Evaluation Metrics
We followed the original evaluation metrics used in the respective tasks. For classification tasks (Character Guessing, Question Answering and Personality Understanding), we report accuracy using exact match. For sequence labeling and coreference tasks (Coreference and Role Detection), we use F1 scores as the evaluation metric. For summarization tasks (Text Summarization), we report ROUGE scores to measure the overlap between the generated and reference summaries.
Task Name Task Type Data Source Metrics
TVshow Guessing Character Guessing TV Shows Exact Match
FriendsQA Question Answering TV Shows Exact Match
Screenplay Character Coreference Movies F1 Score
SummScreen Text Summarization TV Shows ROUGE
CSI Role Detection TV Shows F1 Score
PERSONET Personality Understanding Novels Exact Match
Table 1: Overview of benchmark tasks, including their type, data sources, and evaluation metrics.
3.2 Fictional Works Dataset
The existing six benchmark datasets, which are derived from 91 different TV shows, 12 different movies, and 33 classic literature books. We constructed the Fictional Works Dataset by uniformly sampling 200 segments from 40 selected sources, including 11 distinct TV series, 9 movies, and 20 novels, each source provides 5 scripts. Each segment contains an average of 312 words. The format of each segment is: TV Show Title; Scene, where the scene consists of a mixture of dialogue and narration from the script.
3.3 Extended Dataset Analysis
For the six representative character understanding tasks studied in our work, we selected one dataset for each category. However, we also recognize that there are many different benchmark datasets for each type of task. When selecting representative datasets for each category, we focused on choosing those with more diverse sources. Additionally, we supplemented the selection with eight datasets of the same type but derived from different sources in Table 2.
Task Name Task Type Source
Character Mining Character Identification TV Shows
LISCU Character Identification Book Reviews
MovieQA Visual Question Answering Movies
Liter Coreference Character Coreference Novels
HVV Detect Role Detection Movies, News
SYMON Visual Question Answering Movies
CEB Character Representation Novels
Love and Violence Summarization TV Shows
Table 2: Overview of 14 tasks in the character understanding benchmark.
3.4 Model
We pass all passages with the same prompt through GPT4o (gpt-4o), [^7] https://platform.openai.com/docs/models/gpt-4-and-gpt-4-turbo using the OpenAI API. The cost for the tasks in s 4 and s 5 is approximately \$80, while the cost for the task in s 6 is approximately \$40. The character description used in section  5 is generated by GPT-4 (gpt-4-turbo) API. For the prompts and description generated by GPT4 itself, please refer to Appendix C. Due to the inherent characteristic of ChatGPT’s unstable output, even though we had set the temperature to 0, each generation may exhibit slight variations in content. In our experiments, there is a variation of 1% to 3% in the results. To mitigate this, we repeat each experiment three times and report the average scores.
4 Verbatim Memorization Evaluation
We address our first research question: How does GPT utilize and integrate multi-source information to identify the origin of a script passage?
We do this by considering a task that is focused on evaluating the extent to which models rely on verbatim memory to solve problems without requiring significant reasoning.
4.1 Task Definition
Name Source Prediction with Rationale
In this Source Memorization task, we assess the model’s ability to predict the origin of a text passage, investigating the extent to which it relies on memorized pretraining knowledge. The model is given a segment script and we ask it to identify the source while also providing a rationale for its prediction 9. This task includes two experimental settings: (1) with or without multiple-choice options, and (2) with or without the requirement for a rationale explanation, which made four combinations.
Rationale Analysis
We collected and analyzed the rationales generated by GPT, then asked GPT to performing clustering on all rationales and find groups of rationales with similar pattern. This analysis helps identify common elements the model relies on during source prediction, especially when character names are explicitly mentioned as part of the reasoning.
Effect of Input Length
To explore the impact of input length, we limited the length of text segments provided to the model. We compared its performance in two settings: one with character names and specific mentions provided, and the other with them masked. We hypothesize that verbatim memory plays a lesser role in such tasks, aligning with the human prediction process Brainerd and Reyna (1992).
4.2 Result
Source Prediction
The results presented in Table 3 reveal several notable trends across the different experimental setups. First, the model demonstrates a strong memorization ability when presented with the original script, achieving a consistent 96% accuracy even without any options and under the Chain-of-Thought (CoT) condition. This suggests a significant reliance on memorized content, as the model can accurately retrieve the source purely based on internal knowledge. Furthermore, the use of CoT reasoning substantially enhances the model’s performance across all conditions, improving the retrieval accuracy from 96% to 100% in the case of the original script. This indicates that CoT prompts help guide the model toward more accurate predictions by encouraging a more structured reasoning process. However, when character names are masked (using placeholders like P0, P1), the model’s accuracy decreases, highlighting the importance of specific names as primary retrieval cues. The drop in performance is even more pronounced when names are replaced with generic alternatives, suggesting that the model’s retrieval capability is highly dependent on the exact memorization of character names rather than broader contextual information. This is evidenced by the reduction in accuracy from 96% with the original script to 74% with replaced names. In contrast, replacing notable mentions (such as unique terms or identifiable elements other than character names) has a minimal effect on performance, with the model maintaining a high accuracy of 96%-98%. This further underscores that the model’s retrieval strategy is heavily focused on character names, and it places less emphasis on other narrative details when determining the text’s origin. Overall, these findings highlight the model’s dependency on character-specific memorization, revealing both the strengths and vulnerabilities of its retrieval mechanisms when faced with modifications to key identifying elements.
Condition With Options CoT No Options CoT
Origin Script 96% 100% 96% 98%
Name Masked (P0, P1) 92% 98% 82% 86%
Name Replaced 84% 98% 74% 80%
Notable Mentions Replaced 94% 98% 96% 96%
Name + Mention Replaced 84% 96% 72% 80%
Table 3: We find that CoT consistently improves the model’s prediction accuracy. Character names contribute significantly to accuracy, while notable mentions have a very small impact, likely due to the correlation between their influence and their frequency of occurrence in the text.
Rationale Statistics
We analyzed the 7 rationale content generated by the model across these conditions. The following table (Table 4) shows the frequency of the key rationale elements identified in the model’s output. The most frequently mentioned rationale elements include character name, content of dialogue, and setting, indicating that these factors play a significant role in the model’s reasoning process.
Rationale Element Percentage (%)
Character Name 84%
Notable Mentions 28%
Content of Dialogue 66%
Tone 40%
Humor Style 46%
Setting 46%
Social Dynamics 36%
Table 4: We found that the model primarily relies on the combinations of character names mentioned in the text, especially the main characters, when guessing the source of the text. Although notable mentions are rarely utilized, some distinctive catchphrases, special place names, and event titles are still noticed and leveraged by the model.
Word Limitation
Table 5 shows the model’s source identification accuracy with varying token limits, comparing the original script and the script with replaced names. The model consistently achieves 96% accuracy with the original script, demonstrating strong memorization even with limited context (as few as 25 tokens). In contrast, replacing character names leads to a significant drop in accuracy, especially with shorter passages. For the 25-token input, accuracy falls to 26%, highlighting the model’s heavy reliance on specific names for retrieval. Increasing the token limit to 50 tokens improves accuracy to 48%, but extending to 100 tokens shows no further gains. On average, across all 342 samples, the accuracy for replaced names recovers to 74%, suggesting some benefit from longer contexts but still falling short of the original script’s performance. This indicates that character names play a crucial role in the model’s retrieval process, with limited capacity to compensate through broader context when explicit names are obscured.
Word Limit Original Script Replaced Names
25 Words 96% 26%
50 Words 96% 48%
100 Words 96% 52%
All (Avg. 342) 96% 74%
Table 5: By leveraging verbatim memory, GPT demonstrates a surprisingly high memorization ability even with as few as 25 words. The slower improvement as the number of words increases suggests that source prediction predominantly relies on verbatim memory.
4.3 Analysis
Using text excerpts to prompt the model to guess their source highlights the model’s reliance on verbatim memory. Experiments with word count limitations revealed that the number of words had little impact on performance, further confirming that the model relies on verbatim memory for this task. By employing CoT, we encouraged the model to use gist memory for reasoning, resulting in improved performance and greater robustness against perturbations. This, in turn, provides additional evidence that the model predominantly relies on verbatim memory to solve such tasks.
5 Gist Memory Evaluation
The experiment in Section s 4 demonstrated that applying Chain-of-Thought (CoT) can prompt LLMs to utilize gist memory. In this section, we aim to explore additional methods to encourage LLMs to leverage gist memory effectively.We hypothesize that by reducing the confidence LLMs place in their verbatim memory, they are more likely to rely on gist memory while simultaneously applying reasoning. Building on previous findings that changing character names can significantly degrade performance, we intuitively believe that altering names allows the model to recognize changes in the information embedded within names, such as nationality and gender. Additionally, we pose our second question: Does name replacement affect LLMs’ reliance on gist memory, and if so, to what extent? Moreover, does the choice of name replacement method make a difference?
5.1 Task Definition
We adopt the character guessing  Sang et al. (2022b) format, and we masked the character names mentioned in the script as well. For the prompts please refer to Appendix D.
Name Replacement Strategies
We employed the following name replacement methods:
  • Cross-Language Replacement: Replacing names with uncommon names from other languages, swapping the gender in some cases (GS).
  • Same-Language Replacement: Replacing names with common names from the same language, including gender-swapped (GS) variants.
  • Anonymized Replacement: Replacing all character names with neutral labels (e.g., "Speaker 0").
  • Gender-Swap Only (GS Only): Swapping the gender of the original names without changing the language.
  • Control Tests: Introducing incorrect show names or names from other TV shows to evaluate the model’s susceptibility to data contamination (an example in  11).
5.2 Result
Our analysis focuses on understanding how different replacement strategies affect the model’s predictions. Gender-swapping and cross-cultural name replacements may disrupt the model’s memorized associations more effectively than simple anonymization. This helps us gauge the model’s inference capabilities when it cannot rely solely on pretraining knowledge. Table 6 summarizes the accuracy of GPT-4o across different replacement strategies with and without character descriptions. The "Name Return" task shows that GPT-4o achieves a perfect 100% accuracy, indicating strong memorization of original character names. However, the accuracy significantly drops for other replacement strategies, especially for cross-language and gender-swapped (GS) replacements, where the performance decreases to as low as 20.2% and 15.5%, respectively. Interestingly, anonymization (replacing all names with neutral labels like "Speaker 0" and "Speaker 1") shows a moderate increase in accuracy when descriptions are provided, suggesting the model may rely more heavily on contextual information in these cases. The results indicate that providing character descriptions can slightly improve the model’s performance, particularly in the cases of anonymized and incorrect show names replacements.
Replacement Strategy Without Description With Description
Name Return 100 100
No Replacement 52.1 47.9
Cross-Language 35.5 28.2
Cross-Language (GS) 20.2 21.8
Same Language 34.2 24.1
Same Language (GS) 15.5 16.5
Anonymized 16.7 26.4
GS Only 7.3 8.9
Wrong show names 41.5 43.6
Names from other shows 12.3 31.1
Table 6: The impact of different name replacement methods on GPT-4o’s verbatim memory ability was demonstrated. With reduced confidence in its verbatim memorization, GPT-4o is more likely to rely on reasoning. By providing descriptions, GPT-4o is more inclined to access gist memory and then apply reasoning.
5.3 Analysis
Providing character summarization evaluates two key aspects of GPT’s performance. Without summarization, the task primarily tests verbatim memory, where GPT tends to struggle. However, when character descriptions are provided, the task evaluates both memorization and inference abilities. We observe that as the task complexity increases, GPT relies more on reasoning and less on simple memorization.
6 Reasoning Validation Task
The character guessing task format is quite similar to verbatim memorization evaluation tasks, such as the name cloze task Chang et al. (2023b). We have demonstrated that tasks heavily reliant on verbatim memory can measure its reliance and help transfer the model’s usage to gist memory. Intuitively, most types of character understanding tasks should rely more on gist memory. This leads us to our third question: Can we encourage GPT to rely more on gist memory in different character understanding tasks?
6.1 Benchmark Tasks Description
Generally, character understanding tasks require the model to predict a character’s information based on a given segment of text. For script-based character understanding, the provided texts often include conversations from TV show and movie scripts, while novels typically include a mix of narrative descriptions and dialogue. The six selected benchmarks cover a range of character-centric tasks, designed to test different aspects of character understanding(Table 1):
Character Guessing This task involves identifying the speaker for each utterance in a script. Each utterance within a scene is segmented and input into the model, with the speaker’s name preceding each utterance masked and replaced by a unique special token. Consistent special tokens are used to represent the same speaker within a scene. The objective of the character guessing task is to accurately predict the identity of the speaker associated with each special token.
Coreference Resolution Given a conversation in scripts that contains multiple utterances and n character mention entity c1, c2, …, cn within it, the objective of the coreference resolution task is to assemble all mention entities that refer to the same character in a cluster.
Personality Understanding Given a text snippet describing a character from classic literature along with preceding context, the objective of the personality understanding task is to predict the most likely personality trait depicted in the snippet. The task is formulated as a multi-choice problem with distractor traits, testing the model’s ability to infer character personalities from descriptive text.
Role Detection Given dialogues and scene descriptions from scripts, the objective is to determine whether each token refers to the perpetrator. This task is framed as a multi-label classification problem, where each label corresponds to a character in the script, with a focus on detecting roles within a crime narrative context.
Open Domain Question Answering The input is a dialogue excerpt from the script and a manually crafted question, the objective of the open-domain question-answering task is to extract the specific span in the dialogue that answers the question. This can be viewed as an answer span selection task, where the model is expected to identify spans within the dialogue that contain the answer to the question.
Summarization Given dialogue utterances and scene descriptions from transcripts, the objective of the summarization task is to generate a coherent summary that captures key plot points while excluding irrelevant details.
6.2 Perturbed Method
We implemented the name-changing strategy of swapping genders across cultures, which, as demonstrated in Section s 5, significantly drops performance.
6.3 Result
Table 7 presents the impact of replacing character names on a variety of downstream tasks using the GPT-4o model. The experiments assess performance across six tasks:
Memory-Driven Tasks:
These tasks rely heavily on verbatim memory and are significantly affected by name replacement.
Intuively, TV Show Guessing and Screenplay Coreference should rely more on verbatim memory and they drop of 14.6% and 8.7% meets our expectation.
Reasoning-Driven Tasks:
These tasks primarily rely on gist memory and are more likely to engage in reasoning without external hints(CoT or extra description).
The slight drops in FriendsQA and Personet personality understanding tasks, with a decrease of 0.7% and 2%, align with our expectations. The 17.2% drop in role extraction is unexpectedly high, resembling the trend seen in coreference resolution tasks. The drop in ScreenSum is moderate, falling within expected ranges.
Figure 2: All tasks exhibited performance degradation. The decline in QA and personality understanding tasks was minimal, indicating that GPT effectively leveraged its reasoning abilities for these tasks. The most significant drop was observed in role detection, exceeding our expectations, which suggests that verbatim memory plays a substantial role in supporting this type of task.
Task Name Origin Script Name Replaced
TVShow Guessing 44.8 30.2
Screenplay Coreference 58.3 49.6
FriendsQA 29.3 28.6
ScreenSum (ROUGE_L) 19.1 13.8
CSI Role Detection 50.1 32.9
PERSONET Personality 40.0 38.0
Table 7: All tasks showed a decline in performance, with the smallest drop in QA, which exceeded our expectations. We believe this is because the questions asked were relatively simple and basic. There was almost no change in personality understanding, indicating that the model knows it should use gist memory for such tasks. The largest decline was observed in role extraction, which also exceeded our expectations, suggesting that the model heavily relies on verbatim memory when handling this type of task.
6.4 Analysis
Our experiment, through name replacement, increases the difficulty of the model relying on verbatim memory. As a result, the model is more likely to focus on gist memory, thereby enabling reasoning. By comparing the performance before and after name replacement, we can observe whether the model relies more on verbatim or gist memory without external hints. Specifically, if the performance difference is small, it suggests that the model is more likely to use reasoning for answering such tasks. Conversely, if the difference is large, it indicates that the model tends to rely more on verbatim memory. All tasks showed a decline, and we observed that the model always tends to prioritize verbatim memory. This ’lazy’ behavior suggests that to unlock the model’s true reasoning ability, we need to design methods that encourage the model to focus more on gist memory, such as reducing the similarity between the text and pretraining data.
7 Conclusion
In this paper, we evaluate the extent of GPT-4’s verbatim memory for fictional works and its impact on downstream tasks. We also propose methods to mitigate this impact and encourage the model to use reasoning. We draw several key conclusions: (1) The model always tends to first rely on verbatim memory when performing character understanding tasks, and it consistently benefits from this approach. (2) When the predicted text differs significantly from the pretraining data, the model is more likely to use reasoning to attempt to compensate.
8 Limitations
The specific training data of GPT-4o is not publicly disclosed by OpenAI, making it fundamentally unknowable. Our work employs probabilistic inference to estimate the model’s familiarity with a set of fictional works. However, whether these works are explicitly included in the training data remains unanswerable. Nevertheless, even mere familiarity can suffice to contaminate test data, posing challenges to fair evaluation.
Our study can be extended by future research in a number of ways. First, though we ground and motivate our work through “verbatim” and “gist” memory, we note that any apparent “reasoning” of AI systems may differ fundamentally from those of humans, and we did not pursue human studies to examine this in detail. Further investigation is needed to explore this topic.
Second, the relationship between memorization and reasoning remains a subject of debate in cognitive science. Future studies may uncover insights into this relationship from entirely new perspectives.
Finally, our study was centered on select, recent (but not contemporaneous) majority Western popular culture references. Given their popularity, this selection likely increases the chances that LLMs have been pretrained on more instances or language discussing those references (vs. contemporaneous references, or references from non-Western cultures, where overall occurrences in training data may be lower); it is possible that this exaggerated some of the effects we observed.
9 Ethics
Our study probes the degree to which GPT-4o have memorized fictional works, and its impact on character understanding tasks. This work uses the OpenAI API on experiments above, and at no point do we access, or attempt to access, the true training data behind these models, or any underlying components of the systems. Risks The several datasets and fictional works scripts in our experiment are sourced from publicly available sources. However, we cannot guarantee that they are devoid of socially harmful or toxic language. We use ChatGPT  [^8] https://chatgpt.com/ to correct grammatical errors in this paper.
Appendix A Example of Script-based Character Understanding Task
In Table 8 we show a full example of the character understanding task. In this example, we have automatically replaced the individual character names with placeholders like P0 and P1.
Input P0: Hey, sorry about that P1: No, we’re sorry. We never should have been comparing relationships in the first place. P2: Why? We won. You know, I say, next, we take on Koothrappali and his dog. Really give ourselves a challenge. P3: I just want to say one more thing about this. Just because Penny and I are very different people does not mean that we’re a bad couple. P2: The answer is one simple test away. Hmm? You know, it’s like when I thought there was a possum in my closet. Did I sit around wondering? No, I sent Leonard in with a pointy stick and a bag. P3: I killed his Chewbacca slippers. P0: Let’s just take the test. P3: No, no, no, I don’t want to. P0: Oh, well, ’cause you know we’re gonna do bad. P3: Because it doesn’t matter. I don’t care if we’re a ten or a two. P2: Or a one. A one is possible. P3: Marriage is scary. You’re scared, I’m scared. But it doesn’t make me not want to do it. It, it just makes me want to hold your hand and do it with you. P0: Leonard. P1: It makes me so happy if you said things like that. P2: We got an eight-point-two. Trust me, you’re happy.
Label P0: Penny, P1: Amy, P2: Sheldon, P3: Leonard
Table 8: An example from character guessing task.
Appendix B Example for Verbatim Memory
B.1 Source Prediction
Input Tell me the source for following script:Penny: Hey, sorry about that Amy: No, we’re sorry. We never should have been comparing relationships in the first place. Sheldon: Why? We won. You know, I say, next, we take on Koothrappali and his dog. Really give ourselves a challenge. Leonard: I just want to say one more thing about this. Just because Penny and I are very different people does not mean that we’re a bad couple.
Source The Big Bang Theory
Table 9: An example for source prediction task.
Appendix C Self-generated Character Description
For the self-generated character description, we use following prompts to ask ChatGPT to generate a brief character summarization for each character. The following block-quoted text shows the output we got; importantly, it demonstrates the level of deep knowledge about each of these characters that is baked into the model itself.
Input: Give me a detailed character description for all main characters in the TV show Friends.
Output: Rachel Green (Jennifer Aniston):
Background: Initially introduced as a runaway bride, Rachel starts as a waitress at Central Perk and evolves into a successful fashion executive.
Character Traits: Fashion-conscious, independent, initially somewhat spoiled, but grows into a more self-sufficient and career-focused woman.
Romantic Arc: Has an iconic on-again, off-again relationship with Ross throughout the series.
Ross Geller (David Schwimmer):
Profession: Paleontologist working at a museum.
Character Traits: Intelligent, nerdy, and somewhat socially awkward. Ross has a complicated love life and is often portrayed as the "smart" but emotionally vulnerable member of the group.
Romantic Arc: His tumultuous relationships, particularly with Rachel, are central to the series.
Monica Geller (Courteney Cox):
Profession: Chef and later a restaurant owner.
Character Traits: Competitive, organized, and slightly obsessive about cleanliness. Monica is Ross’s sister and becomes the motherly figure of the group.
Romantic Arc: Her relationship with Chandler Bing becomes a major storyline.
Chandler Bing (Matthew Perry):
Profession: Holds an office job with a mysterious title related to statistical analysis.
Character Traits: Known for his sarcastic humor, commitment issues, and use of humor as a defense mechanism. Evolves into a more mature and emotionally secure character.
Romantic Arc: Develops a significant romantic relationship with Monica.
Joey Tribbiani (Matt LeBlanc):
Profession: Aspiring actor throughout the series.
Character Traits: Easygoing, simple-minded, and often portrayed as the least academically inclined. Known for his catchphrase "How you doin’?" and his love for food.
Romantic Arc: Has numerous short-term relationships, and his genuine friendship with Chandler is a key dynamic.
Phoebe Buffay (Lisa Kudrow):
Profession: Eccentric musician with a unique and mysterious past.
Character Traits: Quirky, free-spirited, and spiritually inclined. Phoebe often provides offbeat humor and memorable songs.
Romantic Arc: Her romantic life includes unique relationships, and her eccentricity adds a distinctive element to the group.
These six characters form the core group of friends in the series, and their individual personalities, relationships, and growth over the ten seasons contribute to the show’s humor, heart, and enduring popularity.
Appendix D Prompts for GPT4
for personality traits extraction ’you are a psychology research assistant designed to help analysis character personalities according to conversations.’ conversation ’Here ends the conversation.Give me a character description for each main character according to this conversation.’
Appendix E Name Replacement and Character Comparison
E.1 Cross-Language Uncommon Name Replacement
With gender swap:[Bojing, Cuixia, Jingjing, Yunsheng, Meilin, Yusong] Without gender swap:[Cuixia, Bojing,Yunsheng,Jingjing,Yusong, Meilin]
Task name Before replacement After replacement
TVShow Monica, Joey, Chandler, Phoebe, Ross, Rachel Bojing, Cuixia, Jingjing, Yunsheng, Meilin, Yusong
CSI Grissom, Catherine, Nick, Sara, Brass Jingbo, Jingwen, Huaqiang, Lihui, Weimin
PERSONET Villefort, The Count, Dantès, Caderousse, Danglars, Albert, Mercédès, Elders, Abbé Busoni, Eugénie, Faria priest, Fernand, Franz, Faria, Morre, Monte Cristo Weina, Nüxia, Danqing, Cailing, Danyu, Ailan, Mengzhe, Lianhua, Busu, Yujun, Fanli, Feiyan, Fangqing, Fangli, Moya, Mingxia
Table 10: Besides the main characters, there are many names of minor characters and passersby appearing in the tasks. In each task, these secondary and passerby names can be dozens of times more than the number of main characters. Replacing these names is complex but not impossible (e.g., with the help of LLMs). However, to align with the TV show approach of only replacing the names of main characters, we only replaced the main characters in all tasks. From the results, it is clear that even replacing only the main characters has had a significant effect.
replace all the character name [Monica, Joey, Chandler, Phoebe, Ross, Rachel]
in the conversation with
[Bojing, Cuixia, Jingjing, Yunsheng, Meilin, Yusong].
The comparison before and after name substitution
The scene overall reflects a group dynamic characteristic of close friends who are comfortable jesting with each other, while also being there in times of crisis or emotional turmoil.
1.Monica: Responsible, caring, and organized. Assertive and confident in her actions.
1. Bojing: Bojing comes across as practical, level-headed and caring. He often acts as the voice of reason for his friends, attempting to mediate, clarify, and console in various situations. His attempts to play down his date suggest he is a private person who doesn’t enjoy sharing intimate details of his life.
2. Joey: Energetic, extroverted, and casual. Lacks the sensitivity of others’ feelings at times but genuinely care about friends.
2. Cuixia: Cuixia seems like a lively, fun-loving character. However, she sometimes shows a more cynical side, quick to suspect something might be wrong with Bojing’s date and suggesting a strip joint as a solution to Meilin’s woes.
3. Chandler: Witty and self-deprecating with an approachable sense of humor. Exhibits insecurity and anxiety in his dialogues, making references to uncomfortable situations and questioning his own actions.
3. Jingjing: Jingjing can be open and candid about his thoughts, even if they seem inappropriate or unusual. He’s also humorously self-aware, undercutting his moments of honesty with reminders that he might be oversharing.
4. Phoebe: Quirky, eccentric, and a free spirit. Her train of thought tends to lean toward the unusual and bizarre. However, she is also compassionate and caring.
4. Yunsheng: Appears offbeat and unusual, suggesting the eating chalk anecdote about his past relationship. He also believes in new-age concepts like auras, showing a more spiritual side.
5. Ross: Insecure and somewhat neurotic and vulnerable. His behavior is indicative of someone who is going through emotional turmoil.
5. Meilin: Exhibits vulnerability and emotional turmoil, especially regarding his recent divorce. He seems to be fluctuating between hurt, anger, and longing for his past relationship.
6. Rachel: Spontaneous and open to change, she takes risks and is adaptable. She is also reliant on her relationships with others, signifying her dependency and need for support.
6. Yusong: Yusong presents as impulsive, high-strung and somewhat comical in moments of panic. Fleeing her wedding because of a sudden realization shows she can make drastic decisions based on her emotions.
E.2 Same-Language Replacement
With Gender-Matched replacement, we replace all the character name [Monica, Joey, Chandler, Phoebe, Ross, Rachel]
in the conversation with
[Sally, Andrew, Ethan, Olivia, Benjamin, Lauren].
With Gender Swap as well, we got [Andrew, Sally,Olivia, Ethan, Lauren,Benjamin].
E.3 Anonymized Speaker Replacement
replace all the character name [Monica, Joey, Chandler, Phoebe, Ross, Rachel]
in the conversation with
[Speaker 0, Speaker 1, Speaker 2, Speaker 3, Speaker 4, Speaker 5]. LLM summarized character description
We use the prompt of "you are a psychology research assistant designed to help analysis character personalities according to conversations……..Here ends the conversation.Give me a character description for each main character according to this conversation." " and we get such result from TVshow Friends scene1.
E.4 Control Test
Example Here is a dialogue between 2 people: Sheldon : (now lifts her head) phoebe!! Amy : you know it’s me? Sheldon : for like a half an hour! man, you can lie about sweden! Amy : how can you come here? Sheldon : how could you not tell me you worked here? Amy : i don’t have to tell you everything!P2: We got an eight-point-two. Trust me, you’re happy.
Prompt The following is a dialogue between the same two people. Please identify who they are:
P0 : (singing) ïpan stripan, glupi glabi! änd that is the swedish national anthem! thank you for asking! (looks annoyed) P1 : wow, ikea… what a rich culture. uhm, you know what? i have a friend who is a masseuse. P0 : oh, ja! ja! P1 : yah! she’s… uhm… not very good though… (phoebe looks devastated)
Table 11: An example for the control test, where in the example,the script is from Friends but speakers are annoted as Amy and Sheldon(from The Big Bang Theory).
Appendix F Verbatim avoid Reasoning
Figure 3: In a personality understanding task where the scenario involves ’Joey returned the lost wallet to its rightful owner, Monica,’ the question posed might be, ’According to this action, his personality should be: ?’ The human reasoning process would draw on contextual evidence, concluding that returning lost property demonstrates honesty. In contrast, the LLM might conclude, ’Joey and Monica are from the TV show Friends, and Joey is characterized as an honest character in this show,’ thus relying heavily on memorized information rather than genuine reasoning.
Appendix G Benchmark Datasets
The six selected benchmarks cover a range of character-centric tasks, designed to test different aspects of character understanding(Table 1):
  • TVShow Guessing: Based on the TVSHOWGUESS dataset, which contains scripts from five popular TV series: Friends, The Big Bang Theory, The Office, Frasier, and Gilmore Girls Sang et al. (2022c), this is a multi-choice classification task. The dialogue is anonymized using placeholder identifiers (e.g., P1, P2), and the model must predict the correct speaker for each line, testing its ability to identify characters based on context.
  • Screenplay: Using screenplays from nine popular movies, such as Avengers: Endgame (2019), Dead Poets Society (1989), John Wick (2014), and The Shawshank Redemption (1994) Baruah and Narayanan (2023), this is a coreference resolution task. The model is tasked with linking character mentions throughout the text, forming coreference clusters and resolving ambiguities in character references.
  • Personality Understanding: Using the PERSONET dataset, which comprises approximately 32,000 samples from 33 classic literature books Yu et al. (2023b), this task involves predicting the most likely personality trait depicted in a given text snippet, along with its preceding context. Formulated as a multi-choice problem with distractor traits, this task evaluates the model’s ability to infer character personalities from descriptive text.
  • CSI Role Detection: Based on 39 episodes of the TV series CSI Frermann et al. (2018), this sequence labeling task requires the model to identify mentions of the perpetrator in dialogues and scene descriptions, focusing on role detection within crime-related narratives.
  • FriendsQA: Derived from 1,222 scenes from the first four seasons of Friends Yang and Choi (2019), this dataset supports an open-domain question answering task. The model is given a dialogue excerpt and a manually crafted question and must extract the specific span in the dialogue that answers the question, testing its comprehension of dialogue context.
  • SUMMSCREEN: Built from transcript-recap pairs of 91 TV series collected from community sources such as The TV MegaSite, Inc. (TMS) and ForeverDreaming (FD) Chen et al. (2022), this dataset supports a summarization task. The model is required to generate a coherent summary from dialogue utterances and scene descriptions, capturing key plot points while excluding irrelevant details.

### v4

@@
 Abstract
-Recently, Large Language Models (LLMs) have shown impressive performance in character understanding tasks, such as analyzing the roles, personalities, and relationships of fictional characters.
+Recently, Large Language Models (LLMs) have shown impressive performance in various character understanding tasks.
 However, the extensive pre-training corpora used by LLMs raise concerns that they may rely on memorizing popular fictional works rather than genuinely understanding and reasoning about them.
-In this work, we argue that “gist memory”—capturing essential meaning—should be the primary mechanism for character understanding tasks, as opposed to “verbatim memory”—exact match of a string.
-We introduce a simple yet effective method to mitigate mechanized memorization in character understanding evaluations while preserving the essential implicit cues needed for comprehension and reasoning.
-Our approach reduces memorization-driven performance on popular fictional works from 96% accuracy to 72% and results in up to an 18% drop in accuracy across various character understanding tasks.
-These findings underscore the issue of data contamination in existing benchmarks, which often measure memorization rather than true character understanding.
+Prior work has addressed memorization in LLM reasoning tasks, but these approaches do not directly apply to character understanding.
+To address this, we leverage the cognitive science concepts of "gist memory" and "verbatim memory", where verbatim memory corresponds to exact recall, while gist memory captures underlying meaning.
+We design a hard-soft intervention framework, where the hard setting disrupts memorization by perturbing key entities, while the soft setting maintains the original text but guides the model toward reasoning.
+Our experiments across six character understanding tasks with GPT-4o, LLaMA3.3-70B, and DeepSeek V3 demonstrate that reducing verbatim memory reliance shifts model behavior toward reasoning.
+Our method leads to a performance drop of up to 45.8% when models can no longer depend on memorization, highlighting the extent to which existing benchmarks inflate performance through data contamination.
+Additionally, we show that prompting techniques can systematically alter a model’s response strategy, making it more dependent on either memorization or reasoning.
 Memorization Over Reasoning? Exposing and Mitigating Verbatim Memorization in Large Language Models’ Character Understanding Evaluation
@@
 By relying on gist memory, the model is encouraged to engage in reasoning rather than purely relying on memorization.
-Motivated by these concepts, we propose an approach that alters elements such as character names, place names, and specific notations in the sentences while preserving the relationships between characters, significant events, and other information we want the model to utilize.
-We hypothesize that these modifications can reduce the model’s tendency to take shortcuts by relying on verbatim memory, and instead encourages the model to rely on gist memory to stimulate its reasoning ability for character understanding tasks.
-We tested our method on six key types of character understanding tasks and obtained an interesting finding: When the difficulty of using verbatim memory increases, the model is more likely to resort to reasoning to compensate.
-Our work provides a simple approach to ensure a fairer benchmark for character understanding tasks on large models that already have strong memorization abilities of fictional works.
-It offers an alternative method compared to more costly solutions, such as developing data from obscure or very new books that the model has not encountered Ngo et al.
-(2024), or using techniques like machine unlearning to forget original training texts Shaik et al.
-(2024).
-In summary, we provide an outline of the paper and summarize our key contributions:
-  • To determine whether the memory types used in character understanding tasks align with our expectations, we draw on human cognitive science to define gist and verbatim memory in the context of character understanding tasks (s 2).
-  • We benchmarked the mechanism memorization ability of over 100 sources of fictional works (s 4.1).
-  • We measured the verbatim memory of large models for hundreds of different source fiction works and tested how different keywords affect the verbatim memory of these models (s 4).
-  • We demonstrated that even for the same factor, such as a person’s name, different disturbances, such as gender or cultural interference, can lead to significantly different verbatim memory effects in large models (s 5).
-  • We compared the different effects of the model using verbatim and gist memory across six different test sets and found that the model’s accuracy decreased by 0.7% to 17.2% (s 6).
-2 Related Task
-Among Large Language Models (LLMs)’ various capabilities Li et al.
-(2024b, d); Zhang et al.
-(2024); Li et al.
-(2024c), the reasoning abilities Tong et al.
-(2023); Yang et al.
-(2023); Xiong et al.
-(2024b); Tong et al.
-(2024); Xiong et al.
-(2024a) of LLMs Li et al.
-(2024a); Tan et al.
-(2024) have garnered significant attention, but their memory has also been shown to play a crucial role in influencing benchmark reasoning performance Zhou et al.
-(2023).
-This influence is complex: data contamination can inflate performance metrics Khandelwal et al.
-(2020); Blevins and Zettlemoyer (2022); Carlini et al.
-(2023), while knowledge conflict can reduce them Su et al.
-(2024); Wang et al.
-(2024); Xie et al.
-(2024b); Xu et al.
-(2024).
-Recent efforts have aimed to mitigate the impact of memory on reasoning benchmarks to establish more reliable evaluations, particularly for reasoning tasks like numerical reasoning, commonsense reasoning, or logical reasoning, where memory requirements are less significant and easier to define Xie et al.
-(2024a); Sainz et al.
-(2023); Li et al.
-(2024e).
-However, in character understanding tasks, the diversity of task formats and scenarios means that memory requirements vary significantly, making it challenging to establish a universal standard to distinguish between acceptable and unacceptable use of memory.
-We propose leveraging gist memory to reduce LLMs’ reliance on precise memorization.
-This approach encourages models to utilize reasoning capabilities more effectively in solving character understanding tasks, aligning them more closely with the original intent behind these benchmarks.
-Character Understanding
+Motivated by the concept of gist memory, we propose a prompt-based approach to mitigate model reliance on verbatim memorization in character understanding tasks.
+We evaluate our method on GPT-4o, LLaMA3-70B, and DeepSeek V3, introducing a hard-soft setting framework: the hard setting blocks memorization by replacing character names, while the soft setting retains the original text but employs gist-based prompts to encourage reasoning over recall.
+Our ablation study confirms the effectiveness of both interventions.
+Building on these findings, we extend our approach to five character understanding tasks, designing task-specific gist prompts and evaluating models under both settings.
+Our results indicate that the Gist+NR approach best reduces memorization dependency.
+In summary, we outline our study and key contributions as follows:
+  1. 1.
+    Character understanding tasks heavily rely on verbatim memorization.
+Across six benchmark tasks, all evaluated LLMs show significant performance drops when explicit memorization is disrupted, confirming that these tasks often assess recall rather than true reasoning (Table 6.2).
+  2. 2.
+    Gist-based prompting effectively reduces memorization dependency.
+Introducing gist-based prompts systematically lowers reliance on verbatim recall across all tasks, encouraging models to use reasoning instead of mechanical memorization (Section 6).
+  3. 3.
+    Verbatim and gist-based prompting modulate model response strategies.
+Models adjust their response strategies based on the prompting style—verbatim-based prompts reinforce reliance on memorization, while gist-based prompts encourage reasoning-driven answers (Section 4.3).
+2 Related Work
+2.1 Character Understanding and its evolution with LLM
 Character understanding has made significant contributions to advancing story comprehension in psychological, literary, and educational research  Bower and Morrow (1990); McKee (2010); Paris and Paris (2003).
@@
 (2023b).
-Recent studies have shown promising performance by large language models (LLMs) on character understanding tasks Li et al.
+Recent studies have shown promising performance by LLMs on character understanding tasks Li et al.
 (2023); Stammbach et al.
@@
 (2023a), are prone to strong memorization.
-In our work, we measured the verbatim capabilities of LLMs on these sources, as well as the potential inflation of benchmark results caused by overly strong verbatim memory.
-Our approach also mitigates this effect.
-Verbatim and Gist Memory
-Cognitive science defines two types of memory: verbatim memory and gist memory Brainerd and Reyna (2002); Reyna and Brainerd (1998).
-Verbatim memory recalls the exact wording of a sentence, such as "The black-and-white cat sat on a soft, red mat in the living room," retaining every detail about the cat’s color, the mat’s texture and color, and the location.
-In contrast, gist memory captures the essential meaning, like "A cat was sitting indoors," focusing on the core idea—the presence of a cat and its general location—while discarding details such as the cat’s color or the specific description of the mat.
-This highlights how verbatim memory retains all specific details, while gist memory abstracts the key information by ignoring less critical details.
-Cognitive science research suggests that humans primarily rely on gist memory for reasoning, and these two types of memory are independent, rather than hierarchical Brainerd and Reyna (1992).
-In recent studies on LLM memorization Carlini et al.
-(2023), the extraction of strings from a larger text is defined as a form of verbatim memory.
-Their solution primarily focuses on generating entirely new test samples that the model has never encountered Xie et al.
+Naturally, this raises curiosity about how memorization affects downstream tasks, particularly character understanding.
+While previous research has demonstrated the memorization capabilities of LLMs, our work further explores its specific impact on character understanding and proposes an approach to mitigate this effect.
+2.2 LLMs’ Memorization and Reasoning
+Among LLMs’ various capabilities Li et al.
+(2024c, e); Zhang et al.
+(2024); Li et al.
+(2024d), the reasoning abilities Tong et al.
+(2023); Yang et al.
+(2023); Xiong et al.
+(2024b); Tong et al.
+(2024); Xiong et al.
+(2024a) of LLMs Li et al.
+(2024b); Tan et al.
+(2024) have garnered significant attention, but their memorization has also been shown to play a crucial role in influencing benchmark reasoning performance Zhou et al.
+(2023); Xie et al.
+(2024a); Chen et al.
+(2024); Li et al.
+(2024a); Biderman et al.
+(2024).
+The influence of memorization is complex: data contamination can inflate performance metrics Khandelwal et al.
+(2020); Blevins and Zettlemoyer (2022); Carlini et al.
+(2023), while knowledge conflict can reduce them Su et al.
+(2024); Wang et al.
+(2024); Xie et al.
+(2024b); Xu et al.
+(2024).
+While prior work has focused on detecting and mitigating memorization effects in reasoning tasks such as numerical, commonsense, and logical reasoning Xie et al.
+(2024a); Sainz et al.
+(2023); Li et al.
+(2024f), character understanding tasks introduce additional challenges due to their diverse formats and varying memory demands.
+Unlike structured reasoning tasks where excessive memorization is clearly undesirable, character understanding often requires a nuanced balance between recall and inference.
+2.3 Verbatim and Gist Memory in LLM Reasoning
+Cognitive science distinguishes between verbatim memory, which retains exact details, and gist memory, which captures the essential meaning Brainerd and Reyna (2002); Reyna and Brainerd (1998).
+For example, verbatim memory would store a sentence like "The black-and-white cat sat on a soft, red mat in the living room" with all details intact, while gist memory would reduce it to "A cat was sitting indoors," focusing on the core idea while discarding specifics.
+Human reasoning primarily relies on gist memory, as the two types of memory function independently rather than hierarchically Brainerd and Reyna (1992).
+In LLM research, memorization is often equated with verbatim memory, where models retrieve and reproduce exact strings from pretraining data Carlini et al.
+(2023).
+A common mitigation strategy involves curating test samples that the model has never seen before Xie et al.
 (2024a).
-However, for character understanding tasks, using obscure or newly released datasets that the model hasn’t seen comes with high costs, and the quality and acquisition costs of the texts are difficult to control.
-Our work demonstrates that using gist memory provides a simple and effective way to alleviate this issue.
+However, in character understanding tasks, acquiring obscure or newly released datasets can be costly and difficult to control in terms of quality.
+Instead, our work proposes leveraging gist memory to reduce reliance on precise memorization, encouraging LLMs to generalize character understanding beyond string-matching.
+By shifting the focus toward gist memory, we aim to align LLM character analysis with human-like reasoning rather than dataset recall.
 3 Data and Experiment Setup
-In our experiments, we utilized two distinct datasets: (1) Benchmark Datasets: six pre-existing datasets designed to evaluate various character understanding tasks, and (2) Fictional Works Dataset: a newly constructed dataset consisting of script excerpts sampled directly from the original texts used to create the benchmark datasets.
-This dataset, drawn from 40 diverse sources—including TV shows, novels, and movies—was specifically designed to investigate GPT’s familiarity with the original texts underlying the benchmark datasets.
-3.1 Benchmark Datasets
-We evaluated the interplay between memory and reasoning across six types of character understanding tasks.
-The datasets used are detailed in this section, while task descriptions are provided in Section 6.1.
+In our work, we utilize two distinct datasets to evaluate LLMs on: (1) their reasoning capabilities in character understanding under gist memory and (2) their memorization of original source texts.
+3.1 Benchmark Datasets for Character Understanding Tasks
+To evaluate the interplay between memorization and reasoning in character understanding tasks, we employ six publicly available datasets covering diverse task types.
+These datasets span different narrative mediums, including TV shows, movies, and novels, ensuring diversity in both content and character representation.
+The task formulations and dataset compositions are detailed below, with an overview provided in Table 1.
+Further task details are available in  6.1.
 Datasets
-For coreference resolution , we use the latest released version of the MovieCoref dataset  [^1] https://github.com/usc-sail/mica-character-coref.
-For character guessing, we adopt the TVSHOWGUESS dataset  [^2] https://github.com/YisiSang/TVSHOWGUESS to conduct experiments.
-For personality understanding, we utilize the PERSONET dataset  [^3] https://github.com/Gorov/personet_acl23.
-For role detection, we use the CSI dataset  [^4] https://github.com/EdinburghNLP/csi-corpus.
-For open-domain question answering, we conduct experiments using the FriendsQA dataset  [^5] https://github.com/emorynlp/FriendsQA.
-For summarization task, we adopt the SummScreen dataset  [^6] https://github.com/mingdachen/SummScreen.
-We follow all the data preprocessing, development, and testing separation provided by the original datasets.
-Tasks and datasets details are in appendix G.
+We evaluate character understanding using six publicly available datasets spanning diverse sources, including TV shows, movies, and novels.
+For coreference resolution, we use the latest release of the MovieCoref dataset  [^1] https://github.com/usc-sail/mica-character-coref, which provides annotated coreference chains in movie scripts.
+For character guess, we adopt TVSHOWGUESS  [^2] https://github.com/YisiSang/TVSHOWGUESS, where models must identify the speaker of a given dialogue.
+For personality understanding, we use PERSONET  [^3] https://github.com/Gorov/personet_acl23, a dataset derived from literary sources with personality trait annotations.
+For role detection, we employ CSI  [^4] https://github.com/EdinburghNLP/csi-corpus, which focuses on identifying key characters in crime narratives.
+For open-domain question answering, we evaluate models on FriendsQA  [^5] https://github.com/emorynlp/FriendsQA, where questions are answered by extracting spans from TV show dialogues.
+Lastly, for summarization, we use SummScreen  [^6] https://github.com/mingdachen/SummScreen, which pairs TV show transcripts with corresponding summaries.
+We follow the original dataset authors’ preprocessing, development, and test splits to ensure comparability.
+Further dataset details are provided in Appendix G.
 Evaluation Metrics
-We followed the original evaluation metrics used in the respective tasks.
-For classification tasks (Character Guessing, Question Answering and Personality Understanding), we report accuracy using exact match.
-For sequence labeling and coreference tasks (Coreference and Role Detection), we use F1 scores as the evaluation metric.
-For summarization tasks (Text Summarization), we report ROUGE scores to measure the overlap between the generated and reference summaries.
+We followed the original evaluation metrics used in the respective tasks, shown in Table 1.
 Task Name Task Type Data Source Metrics
-TVshow Guessing Character Guessing TV Shows Exact Match
+TVShow Guess Character Guessing TV Shows Exact Match
 FriendsQA Question Answering TV Shows Exact Match
@@
 Table 1: Overview of benchmark tasks, including their type, data sources, and evaluation metrics.
-3.2 Fictional Works Dataset
-The existing six benchmark datasets, which are derived from 91 different TV shows, 12 different movies, and 33 classic literature books.
-We constructed the Fictional Works Dataset by uniformly sampling 200 segments from 40 selected sources, including 11 distinct TV series, 9 movies, and 20 novels, each source provides 5 scripts.
-Each segment contains an average of 312 words.
-The format of each segment is: TV Show Title; Scene, where the scene consists of a mixture of dialogue and narration from the script.
-3.3 Extended Dataset Analysis
-For the six representative character understanding tasks studied in our work, we selected one dataset for each category.
-However, we also recognize that there are many different benchmark datasets for each type of task.
-When selecting representative datasets for each category, we focused on choosing those with more diverse sources.
-Additionally, we supplemented the selection with eight datasets of the same type but derived from different sources in Table 2.
-Task Name Task Type Source
-Character Mining Character Identification TV Shows
-LISCU Character Identification Book Reviews
-MovieQA Visual Question Answering Movies
-Liter Coreference Character Coreference Novels
-HVV Detect Role Detection Movies, News
-SYMON Visual Question Answering Movies
-CEB Character Representation Novels
-Love and Violence Summarization TV Shows
-Table 2: Overview of 14 tasks in the character understanding benchmark.
-3.4 Model
-We pass all passages with the same prompt through GPT4o (gpt-4o), [^7] https://platform.openai.com/docs/models/gpt-4-and-gpt-4-turbo using the OpenAI API.
-The cost for the tasks in s 4 and s 5 is approximately \$80, while the cost for the task in s 6 is approximately \$40.
-The character description used in section  5 is generated by GPT-4 (gpt-4-turbo) API.
-For the prompts and description generated by GPT4 itself, please refer to Appendix C.
-Due to the inherent characteristic of ChatGPT’s unstable output, even though we had set the temperature to 0, each generation may exhibit slight variations in content.
-In our experiments, there is a variation of 1% to 3% in the results.
-To mitigate this, we repeat each experiment three times and report the average scores.
-4 Verbatim Memorization Evaluation
-We address our first research question: How does GPT utilize and integrate multi-source information to identify the origin of a script passage?
-We do this by considering a task that is focused on evaluating the extent to which models rely on verbatim memory to solve problems without requiring significant reasoning.
-4.1 Task Definition
-Name Source Prediction with Rationale
-In this Source Memorization task, we assess the model’s ability to predict the origin of a text passage, investigating the extent to which it relies on memorized pretraining knowledge.
-The model is given a segment script and we ask it to identify the source while also providing a rationale for its prediction 9.
-This task includes two experimental settings: (1) with or without multiple-choice options, and (2) with or without the requirement for a rationale explanation, which made four combinations.
-Rationale Analysis
-We collected and analyzed the rationales generated by GPT, then asked GPT to performing clustering on all rationales and find groups of rationales with similar pattern.
-This analysis helps identify common elements the model relies on during source prediction, especially when character names are explicitly mentioned as part of the reasoning.
-Effect of Input Length
-To explore the impact of input length, we limited the length of text segments provided to the model.
-We compared its performance in two settings: one with character names and specific mentions provided, and the other with them masked.
-We hypothesize that verbatim memory plays a lesser role in such tasks, aligning with the human prediction process Brainerd and Reyna (1992).
-4.2 Result
-Source Prediction
-The results presented in Table 3 reveal several notable trends across the different experimental setups.
-First, the model demonstrates a strong memorization ability when presented with the original script, achieving a consistent 96% accuracy even without any options and under the Chain-of-Thought (CoT) condition.
-This suggests a significant reliance on memorized content, as the model can accurately retrieve the source purely based on internal knowledge.
-Furthermore, the use of CoT reasoning substantially enhances the model’s performance across all conditions, improving the retrieval accuracy from 96% to 100% in the case of the original script.
-This indicates that CoT prompts help guide the model toward more accurate predictions by encouraging a more structured reasoning process.
-However, when character names are masked (using placeholders like P0, P1), the model’s accuracy decreases, highlighting the importance of specific names as primary retrieval cues.
-The drop in performance is even more pronounced when names are replaced with generic alternatives, suggesting that the model’s retrieval capability is highly dependent on the exact memorization of character names rather than broader contextual information.
-This is evidenced by the reduction in accuracy from 96% with the original script to 74% with replaced names.
-In contrast, replacing notable mentions (such as unique terms or identifiable elements other than character names) has a minimal effect on performance, with the model maintaining a high accuracy of 96%-98%.
-This further underscores that the model’s retrieval strategy is heavily focused on character names, and it places less emphasis on other narrative details when determining the text’s origin.
-Overall, these findings highlight the model’s dependency on character-specific memorization, revealing both the strengths and vulnerabilities of its retrieval mechanisms when faced with modifications to key identifying elements.
-Condition With Options CoT No Options CoT
-Origin Script 96% 100% 96% 98%
-Name Masked (P0, P1) 92% 98% 82% 86%
-Name Replaced 84% 98% 74% 80%
-Notable Mentions Replaced 94% 98% 96% 96%
-Name + Mention Replaced 84% 96% 72% 80%
-Table 3: We find that CoT consistently improves the model’s prediction accuracy.
-Character names contribute significantly to accuracy, while notable mentions have a very small impact, likely due to the correlation between their influence and their frequency of occurrence in the text.
-Rationale Statistics
-We analyzed the 7 rationale content generated by the model across these conditions.
-The following table (Table 4) shows the frequency of the key rationale elements identified in the model’s output.
-The most frequently mentioned rationale elements include character name, content of dialogue, and setting, indicating that these factors play a significant role in the model’s reasoning process.
-Rationale Element Percentage (%)
-Character Name 84%
-Notable Mentions 28%
-Content of Dialogue 66%
-Tone 40%
-Humor Style 46%
-Setting 46%
-Social Dynamics 36%
-Table 4: We found that the model primarily relies on the combinations of character names mentioned in the text, especially the main characters, when guessing the source of the text.
-Although notable mentions are rarely utilized, some distinctive catchphrases, special place names, and event titles are still noticed and leveraged by the model.
-Word Limitation
-Table 5 shows the model’s source identification accuracy with varying token limits, comparing the original script and the script with replaced names.
-The model consistently achieves 96% accuracy with the original script, demonstrating strong memorization even with limited context (as few as 25 tokens).
-In contrast, replacing character names leads to a significant drop in accuracy, especially with shorter passages.
-For the 25-token input, accuracy falls to 26%, highlighting the model’s heavy reliance on specific names for retrieval.
-Increasing the token limit to 50 tokens improves accuracy to 48%, but extending to 100 tokens shows no further gains.
-On average, across all 342 samples, the accuracy for replaced names recovers to 74%, suggesting some benefit from longer contexts but still falling short of the original script’s performance.
-This indicates that character names play a crucial role in the model’s retrieval process, with limited capacity to compensate through broader context when explicit names are obscured.
-Word Limit Original Script Replaced Names
-25 Words 96% 26%
-50 Words 96% 48%
-100 Words 96% 52%
-All (Avg. 342) 96% 74%
-Table 5: By leveraging verbatim memory, GPT demonstrates a surprisingly high memorization ability even with as few as 25 words.
-The slower improvement as the number of words increases suggests that source prediction predominantly relies on verbatim memory.
-4.3 Analysis
-Using text excerpts to prompt the model to guess their source highlights the model’s reliance on verbatim memory.
-Experiments with word count limitations revealed that the number of words had little impact on performance, further confirming that the model relies on verbatim memory for this task.
-By employing CoT, we encouraged the model to use gist memory for reasoning, resulting in improved performance and greater robustness against perturbations.
-This, in turn, provides additional evidence that the model predominantly relies on verbatim memory to solve such tasks.
-5 Gist Memory Evaluation
-The experiment in Section s 4 demonstrated that applying Chain-of-Thought (CoT) can prompt LLMs to utilize gist memory.
-In this section, we aim to explore additional methods to encourage LLMs to leverage gist memory effectively.We hypothesize that by reducing the confidence LLMs place in their verbatim memory, they are more likely to rely on gist memory while simultaneously applying reasoning.
-Building on previous findings that changing character names can significantly degrade performance, we intuitively believe that altering names allows the model to recognize changes in the information embedded within names, such as nationality and gender.
-Additionally, we pose our second question: Does name replacement affect LLMs’ reliance on gist memory, and if so, to what extent?
-Moreover, does the choice of name replacement method make a difference?
-5.1 Task Definition
-We adopt the character guessing  Sang et al. (2022b) format, and we masked the character names mentioned in the script as well. For the prompts please refer to Appendix D.
+3.2 Fictional Works Memorization Benchmark Dataset
+This dataset is designed to evaluate LLMs’ memorization of fictional scripts by assessing their ability to recall verbatim details from diverse narrative sources.
+Benchmark Sources
+The six benchmark datasets described in Section 6.1 are derived from a diverse set of sources, including 91 TV shows, 12 movies, and 33 classic literary works.
+Dataset Construction
+To ensure broad coverage across different narrative mediums, we sample 200 script segments from 40 selected sources, comprising 11 TV series, 9 movies, and 20 novels.
+Each source contributes 5 excerpts, with an average segment length of 312 words.
+Segment Format
+Each excerpt follows the format: TV Show Title; Scene, where the scene consists of a mixture of dialogue and narration extracted from the original script.
+3.3 Models
+we evaluate both open-source and closed-source models, leveraging their APIs for inference.
+Our evaluation includes three frontier LLMs: the open-source Llama 3.3-70B AI (2024) and DeepSeek V3 Liu et al.
+(2024), as well as the closed-source GPT-4o OpenAI (2023).
+4 Soft Setting: Memorization-Controlled Prompting in Character Guessing
+First, we investigate whether prompting strategies can subtly influence LLMs’ response patterns in a soft setting, where the model is not explicitly restricted from using memorization but is instead encouraged to rely more on reasoning over direct retrieval.
+4.1 Character Guessing Task Format
+Let a dialogue segment be represented as $D=\{U_{1},U_{2},...,U_{n}\}$, where each utterance $U_{i}$ is paired with a speaker label $S_{i}$.
+In the Character Guessing Task, the model is given a dialogue excerpt without speaker annotations and must predict the correct speaker for each utterance, $f_{\text{guess}}(U_{i})\rightarrow S_{i}$.
+We follow the original format and setting Sang et al.
+(2022b), using the TVShow Guess Dataset described in Section 3.1.
+This task is designed to prompt the model to make guesses based on character relationships, key events, and personality traits within the text.
+However, its format closely resembles the name cloze tests Chang et al.
+(2023b) recently used to evaluate LLM memorization.
+4.2 Prompting Strategies
+We design two distinct prompting conditions to examine how LLMs rely on different memory mechanisms when identifying characters.
+Verbatim prompting encourages mechanical memorization, where the model recalls exact substrings from its training data and reconstructs missing parts without reasoning.
+In contrast, gist prompting focuses on extracting key information and performing memory-based matching.
+For character understanding tasks, the most relevant gist cues typically include character relationships, significant events, and personality traits.
+All prompts share the same base instruction: "The following is a dialogue. Please identify who they are. …"
+The two variations are as follows:
+Verbatim Prompt: "…Do this by direct memory retrieval—do not infer based on traits, interactions, or context. Recall names exactly from past dialogue knowledge without explanations or reasoning."
+Gist Prompt: "…Do this by analyzing relationships, key events, and personality traits to match them with known characters. Do not directly retrieve from memory who said the sentences."
+4.3 Result and Analysis
+Figure 2 visualizes the performance of the three models on the Character Guessing task under different prompting conditions.
+The performance trends remain consistent across models, and our key observations are as follows:
+  • Baseline Performance: The baseline condition reflects the overall performance of the models, with GPT-4o demonstrating a clear lead, while Llama 3.3-70B and DeepSeek V3 exhibit similar performance.
+  • Memorization Reliance: All three models show that the Verbatim condition performs nearly identically to the baseline, supporting our hypothesis that models primarily rely on memorization when answering the Character Guessing task.
+  • Effect of Gist Prompting: The Gist condition results in a non-trivial performance drop, suggesting that our prompt effectively reduces the models’ reliance on memorization when generating responses.
+Figure 2: The bar chart compares the performance of three LLMs under three prompting settings.
+Across all three models, the performance of the Verbatim and Baseline conditions is highly similar, confirming our hypothesis that the models rely on memorization when answering questions.
+The Gist condition, however, shows a non-trivial decline in performance, indicating that the models struggle when they cannot fully rely on memorization to generate responses.
+5 Hard Setting: Memorization-Controlled Prompting under Name Replacement
+Furthermore, we aim to validate whether our prompting strategies can genuinely alter the model’s reasoning patterns under a hard setting, where memorization is explicitly restricted.
+Previous research has suggested that name replacement may impact model performance Yu et al.
+(2022).
+In this section, we confirm this hypothesis by demonstrating that name replacement significantly affects the model’s memorization ability and quantifying its impact.
+Then, we examine how prompting strategies influence LLMs’ reliance on memorization versus reasoning when memorization is constrained by name replacement.
+5.1 Name Replacement Strategies and Their Impact on Memorization
 Name Replacement Strategies
-We employed the following name replacement methods:
-  • Cross-Language Replacement: Replacing names with uncommon names from other languages, swapping the gender in some cases (GS).
-  • Same-Language Replacement: Replacing names with common names from the same language, including gender-swapped (GS) variants.
-  • Anonymized Replacement: Replacing all character names with neutral labels (e.g., "Speaker 0").
-  • Gender-Swap Only (GS Only): Swapping the gender of the original names without changing the language.
-  • Control Tests: Introducing incorrect show names or names from other TV shows to evaluate the model’s susceptibility to data contamination (an example in  11).
-5.2 Result
-Our analysis focuses on understanding how different replacement strategies affect the model’s predictions.
-Gender-swapping and cross-cultural name replacements may disrupt the model’s memorized associations more effectively than simple anonymization.
-This helps us gauge the model’s inference capabilities when it cannot rely solely on pretraining knowledge.
-Table 6 summarizes the accuracy of GPT-4o across different replacement strategies with and without character descriptions.
-The "Name Return" task shows that GPT-4o achieves a perfect 100% accuracy, indicating strong memorization of original character names.
-However, the accuracy significantly drops for other replacement strategies, especially for cross-language and gender-swapped (GS) replacements, where the performance decreases to as low as 20.2% and 15.5%, respectively.
-Interestingly, anonymization (replacing all names with neutral labels like "Speaker 0" and "Speaker 1") shows a moderate increase in accuracy when descriptions are provided, suggesting the model may rely more heavily on contextual information in these cases.
-The results indicate that providing character descriptions can slightly improve the model’s performance, particularly in the cases of anonymized and incorrect show names replacements.
-Replacement Strategy Without Description With Description
-Name Return 100 100
-No Replacement 52.1 47.9
-Cross-Language 35.5 28.2
-Cross-Language (GS) 20.2 21.8
-Same Language 34.2 24.1
-Same Language (GS) 15.5 16.5
-Anonymized 16.7 26.4
-GS Only 7.3 8.9
-Wrong show names 41.5 43.6
-Names from other shows 12.3 31.1
-Table 6: The impact of different name replacement methods on GPT-4o’s verbatim memory ability was demonstrated.
-With reduced confidence in its verbatim memorization, GPT-4o is more likely to rely on reasoning.
-By providing descriptions, GPT-4o is more inclined to access gist memory and then apply reasoning.
-5.3 Analysis
-Providing character summarization evaluates two key aspects of GPT’s performance.
-Without summarization, the task primarily tests verbatim memory, where GPT tends to struggle.
-However, when character descriptions are provided, the task evaluates both memorization and inference abilities.
-We observe that as the task complexity increases, GPT relies more on reasoning and less on simple memorization.
-6 Reasoning Validation Task
-The character guessing task format is quite similar to verbatim memorization evaluation tasks, such as the name cloze task Chang et al.
-(2023b).
-We have demonstrated that tasks heavily reliant on verbatim memory can measure its reliance and help transfer the model’s usage to gist memory.
-Intuitively, most types of character understanding tasks should rely more on gist memory.
-This leads us to our third question: Can we encourage GPT to rely more on gist memory in different character understanding tasks?
+Previous studies Yu et al.
+(2022) have primarily employed Same-Cultural Name Replacement, a naive and intuitive approach that substitutes character names with culturally consistent alternatives.
+To systematically analyze the impact of name replacement on model performance, we introduce three distinct levels of name modifications:
+  • Name Masking: Replacing all character names with placeholders (e.g., Sheldon → P0, Amy → P1).
+  • Cross-Cultural Name Replacement: Substituting names with those from a different cultural background while maintaining gender consistency (e.g., SheldonJie Zhang, AmyXiaoling Wang).
+This tests whether models rely on specific name distributions for character inference.
+  • Same-Cultural Name Replacement: Replacing names with other culturally consistent alternatives (e.g., SheldonAndrew, AmySally).
+This retains cultural familiarity while disrupting direct memorization.
+Memorization Evaluation Task Format
+We design the Source Identification Task where the model is given a dialogue excerpt without source annotations and must predict the correct origin of the text.
+Let a dialogue segment be represented as $D=\{U_{1},U_{2},...,U_{n}\}$, where each utterance $U_{i}$ is part of a larger script from a specific source.
+The task requires the model to determine the correct source, $f_{\text{source}}(D)\rightarrow S$, where $S$ denotes the source work.
+We use the Fictional Works Memorization Benchmark Dataset described in Section 3.2.
+Impact of Name Replacement on Memorization.
+According to Figure 3, our results demonstrate that different name replacement strategies have a significant impact on the memorization ability of LLMs.
+Among them, Cross-Cultural Name Replacement consistently induces the strongest forgetting effect across all models, indicating that altering names to those from a different cultural background disrupts memorized associations more effectively than other replacement methods.
+Figure 3: The heatmap presents accuracy percentages for source identification across three LLMs under different name replacement conditions.
+All three models demonstrate strong memorization of the script.
+The Cross-Culture Name Replacement method exhibits a significantly strong and the most pronounced memorization-weakening effect across all three models.
+5.2 Memorization-Controlled Prompting under Name Replacement
+Different name replacement strategies have shown a significant impact on the memorization ability of LLMs.
+Building on this, we apply the name replacement strategies introduced in Section 5.1 to evaluate the Character Guessing task using the different prompting strategies discussed in Section 4.2.
+Under the Name Replacement condition, we provide brief character descriptions generated by GPT-4, aiming to enable the model to complete the Character Guessing task through character matching.
+The descriptions used are provided in the Appendix D.
+To assess the actual effectiveness of these descriptions, we conduct an ablation study.
+Figure 4: The line chart shows the decline in memorization across different settings, illustrating that as memorization weakens, overall model performance also decreases.
+The bar chart presents the effect of different prompting strategies: under Name Replacement conditions, Verbatim prompting consistently outperforms the baseline by reinforcing memorization, whereas Gist prompting further reduces reliance on memorization, leading to the lowest accuracy.
+Results and Key Findings
+We present the results in Figure 4 and summarize our main findings as follows:
+  1. 1.
+    Strong Correlation Between Memorization and Performance. Model performance aligns closely with memorization levels, indicating a strong dependence on memorization for accurate predictions.
+  2. 2.
+    Minimal Impact of Character Descriptions .
+As an ablation study, we evaluate the effect of character descriptions under the Origin Script condition.
+Our results show that incorporating character descriptions does not significantly affect model performance, ruling out potential interference from inaccurate or misleading descriptions.
+  3. 3.
+    Verbatim Prompting Enhances Memorization.
+Across all three name replacement strategies, Verbatim prompting consistently improves model performance over the baseline, demonstrating that prompting can reinforce memorization and consequently boost accuracy.
+  4. 4.
+    Gist Prompting Reduces Reliance on Memorization.
+Performance under the Gist condition declines compared to the baseline, confirming that our Gist prompt effectively reduces the model’s reliance on memorization.
+  5. 5.
+    Name Replacement + Gist Prompting as an Effective Evaluation Strategy.
+We conclude that combining Cross-Culture Name Replacement with Gist prompting provides a reasonable and effective method for evaluating a model’s true capabilities while minimizing reliance on memorization.
+6 Evaluating Character Understanding Tasks Under Memorization Constraints
+In this section, we evaluate different types of character understanding tasks using the Gist Prompting with Cross Culture Name Replacement approach.
+We first introduce the design principles for Gist Prompts across various task types, followed by an analysis of the results obtained from three different LLMs.
 6.1 Benchmark Tasks Description
-Generally, character understanding tasks require the model to predict a character’s information based on a given segment of text.
-For script-based character understanding, the provided texts often include conversations from TV show and movie scripts, while novels typically include a mix of narrative descriptions and dialogue.
-The six selected benchmarks cover a range of character-centric tasks, designed to test different aspects of character understanding(Table 1):
-Character Guessing This task involves identifying the speaker for each utterance in a script.
-Each utterance within a scene is segmented and input into the model, with the speaker’s name preceding each utterance masked and replaced by a unique special token.
-Consistent special tokens are used to represent the same speaker within a scene.
-The objective of the character guessing task is to accurately predict the identity of the speaker associated with each special token.
-Coreference Resolution Given a conversation in scripts that contains multiple utterances and n character mention entity c1, c2, …, cn within it, the objective of the coreference resolution task is to assemble all mention entities that refer to the same character in a cluster.
-Personality Understanding Given a text snippet describing a character from classic literature along with preceding context, the objective of the personality understanding task is to predict the most likely personality trait depicted in the snippet.
-The task is formulated as a multi-choice problem with distractor traits, testing the model’s ability to infer character personalities from descriptive text.
-Role Detection Given dialogues and scene descriptions from scripts, the objective is to determine whether each token refers to the perpetrator.
-This task is framed as a multi-label classification problem, where each label corresponds to a character in the script, with a focus on detecting roles within a crime narrative context.
-Open Domain Question Answering The input is a dialogue excerpt from the script and a manually crafted question, the objective of the open-domain question-answering task is to extract the specific span in the dialogue that answers the question.
-This can be viewed as an answer span selection task, where the model is expected to identify spans within the dialogue that contain the answer to the question.
-Summarization Given dialogue utterances and scene descriptions from transcripts, the objective of the summarization task is to generate a coherent summary that captures key plot points while excluding irrelevant details.
-6.2 Perturbed Method
-We implemented the name-changing strategy of swapping genders across cultures, which, as demonstrated in Section s 5, significantly drops performance.
-6.3 Result
-Table 7 presents the impact of replacing character names on a variety of downstream tasks using the GPT-4o model. The experiments assess performance across six tasks:
-Memory-Driven Tasks:
-These tasks rely heavily on verbatim memory and are significantly affected by name replacement.
-Intuively, TV Show Guessing and Screenplay Coreference should rely more on verbatim memory and they drop of 14.6% and 8.7% meets our expectation.
-Reasoning-Driven Tasks:
-These tasks primarily rely on gist memory and are more likely to engage in reasoning without external hints(CoT or extra description).
-The slight drops in FriendsQA and Personet personality understanding tasks, with a decrease of 0.7% and 2%, align with our expectations.
-The 17.2% drop in role extraction is unexpectedly high, resembling the trend seen in coreference resolution tasks.
-The drop in ScreenSum is moderate, falling within expected ranges.
-Figure 2: All tasks exhibited performance degradation.
-The decline in QA and personality understanding tasks was minimal, indicating that GPT effectively leveraged its reasoning abilities for these tasks.
-The most significant drop was observed in role detection, exceeding our expectations, which suggests that verbatim memory plays a substantial role in supporting this type of task.
-Task Name Origin Script Name Replaced
-TVShow Guessing 44.8 30.2
-Screenplay Coreference 58.3 49.6
-FriendsQA 29.3 28.6
-ScreenSum (ROUGE_L) 19.1 13.8
-CSI Role Detection 50.1 32.9
-PERSONET Personality 40.0 38.0
-Table 7: All tasks showed a decline in performance, with the smallest drop in QA, which exceeded our expectations.
-We believe this is because the questions asked were relatively simple and basic.
-There was almost no change in personality understanding, indicating that the model knows it should use gist memory for such tasks.
-The largest decline was observed in role extraction, which also exceeded our expectations, suggesting that the model heavily relies on verbatim memory when handling this type of task.
-6.4 Analysis
-Our experiment, through name replacement, increases the difficulty of the model relying on verbatim memory.
-As a result, the model is more likely to focus on gist memory, thereby enabling reasoning.
-By comparing the performance before and after name replacement, we can observe whether the model relies more on verbatim or gist memory without external hints.
-Specifically, if the performance difference is small, it suggests that the model is more likely to use reasoning for answering such tasks.
-Conversely, if the difference is large, it indicates that the model tends to rely more on verbatim memory.
-All tasks showed a decline, and we observed that the model always tends to prioritize verbatim memory.
-This ’lazy’ behavior suggests that to unlock the model’s true reasoning ability, we need to design methods that encourage the model to focus more on gist memory, such as reducing the similarity between the text and pretraining data.
+We define the five character understanding tasks being benchmarked and provide an illustrative example.
+Let a dialogue segment be represented as $D=\{U_{1},U_{2},...,U_{n}\}$, where $U_{i}$ is an utterance in the conversation.
+Each utterance $U_{i}$ is associated with a speaker label $S_{i}$, and additional annotations depending on the task.
+  • Coreference Resolution: Cluster character mentions $C=\{c_{1},c_{2},...,c_{n}\}$ into groups referring to the same entity, $f_{\text{coref}}(C)\rightarrow\{\text{clusters of }c_{i}\}$.
+  • Personality Understanding: Classify a character’s personality trait $P$ given dialogue context $D$, $f_{\text{personality}}(D,S_{i})\rightarrow P$.
+  • Role Detection: Determine whether a character plays a key role (e.g., perpetrator in a crime narrative), $f_{\text{role}}(S_{i})\rightarrow\{0,1\}$.
+  • Open-Domain Question Answering: Extract the span $A$ in $D$ that answers a given question $Q$, $f_{\text{QA}}(D,Q)\rightarrow A$.
+  • Summarization: Generate a summary $S$ that captures key events and relationships, $f_{\text{summ}}(D)\rightarrow S$.
+Example Dialogue
+To illustrate these tasks, consider the following conversation:
+P0: "You’re still going to have to convince a jury that I killed two strangers for no reason."
+Grissom doesn’t look worried. He takes his gloves off and puts them on the table.
+P1: "You ever been to the theater, Peter? There’s a play called Six Degrees of Separation."
+The correct speaker labels are P0 → Peter Berglund, P1 → Grissom.
+For all tasks except Character Guess, the placeholders P0 and P1 are replaced with their original names, Peter Berglund and Grissom.
+Table 2 illustrates how this dialogue segment is utilized across different tasks.
+Task Example Question and Expected Output
+Coreference Resolution "Which entity does ’his’ refer to?" → "Grissom"
+Personality Understanding "What trait best describes P1?" → Analytical
+Role Detection "Identify all entities referring to the criminal." → Peter Berglund
+Question Answering "What play does Grissom mention?" → "Six Degrees of Separation"
+Summarization "Generate a summary." → "Peter discusses convincing a jury, while Grissom references a play."
+Table 2: Examples of different character understanding tasks applied to the same dialogue.
+6.2 Gist Prompts Across Character Understanding Tasks
+General Guidelines for Gist Prompting
+  • Encourage Structural Reasoning: Guide models to infer character identity and relationships from discourse patterns rather than relying on explicit name associations.
+  • Eliminate Memorization Bias: Ensure prompts emphasize evidence-based inference rather than knowledge retrieval.
+  • Preserve Narrative Coherence: In summarization and question answering, focus on maintaining logical flow rather than phrase-matching.
+Detailed gist prompts are included in Appendix H.
+{NiceTabular}
+l|ccc|ccc|ccc Tasks GPT-4o LLaMA3.3-70B DeepSeek V3
+Origin NR NR+GIST Origin NR NR+GIST Origin NR NR+GIST
+Tvshow Guess 78.2 32.0 $\downarrow$46.2 31.5 $\downarrow$0.5 61.4 31.9 $\downarrow$29.5 27.8 $\downarrow$4.1 60.8 29.7 $\downarrow$31.1 26.7 $\downarrow$3.0
+Coreference 58.3 49.6 $\downarrow$8.7 47.5 $\downarrow$2.1 56.7 48.3 $\downarrow$8.4 48.1 $\downarrow$0.2 47.4 34.6 $\downarrow$12.8 32.3 $\downarrow$2.3
+FriendsQA 44.3 38.6 $\downarrow$5.7 38.6 $\uparrow$0.0 45.7 42.9 $\downarrow$2.8 44.3 $\uparrow$1.4 42.8 39.4 $\downarrow$3.4 34.3 $\downarrow$5.1
+ScreenSum 37.5 28.7 $\downarrow$8.8 23.1 $\downarrow$5.6 40.1 25.9 $\downarrow$14.2 21.3 $\downarrow$4.6 35.6 19.3 $\downarrow$16.3 17.1 $\downarrow$2.2
+CSI Role Extract 50.1 32.9 $\downarrow$17.2 31.7 $\downarrow$1.2 48.3 34.7 $\downarrow$13.6 32.7 $\downarrow$2.0 45.1 29.8 $\downarrow$15.3 23.7 $\downarrow$6.1
+PERSONET 56.0 52.0 $\downarrow$4.0 54.0 $\uparrow$2.0 62.0 46.0 $\downarrow$16.0 44.0 $\downarrow$2.0 44.0 40.0 $\downarrow$4.0 38.0 $\downarrow$2.0
+Table 3: Ablation study examining the impact of name replacement and the GIST method on model performance across six NLP tasks.The subscripted values in the NR column indicate the performance drop relative to the original script, while the subscripted values in the NR+GIST column show the performance change compared to the Name Replace condition.
+6.3 Results and Key Findings
+Our main results and findings are summarized as follows:
+  1. 1.
+    Significant Performance Drop with Name Replacement: Across all tasks, we observe a substantial decline in performance when applying the name replacement (NR) method in Table 6.2.
+This result indicates that reducing explicit memorization has a strong impact on model performance, confirming our hypothesis that existing character understanding tasks rely heavily on memorization.
+Moreover, it highlights the potential inflation of benchmark scores under standard evaluation settings.
+  2. 2.
+    Further Degradation with Gist Prompting: The additional performance drop induced by Gist prompting demonstrates its effectiveness across various character understanding tasks.
+This suggests that Gist prompting can further encourage models to rely less on memorization, reinforcing our findings on the role of memorization in these tasks.
+  3. 3.
+    Memorization Can Be Detrimental to Character Understanding: Interestingly, we observe performance gains in certain cases under the combined NR+Gist setting.
+Our error case analysis reveals that these improvements stem from resolving knowledge conflicts—instances where the original memorization-based responses misled the model, whereas reducing memorization mitigated such errors, leading to better predictions.
+  4. 4.
+    Consistent Trends Across Models: As shown in Figure 5, despite variations in absolute performance, different models exhibit similar trends in performance degradation across all tasks.
+This consistency suggests that various task designs interact similarly with memorization, reinforcing the broader implications of our findings.
+Figure 5: Performance degradation across LLMs due to the effects of GIST+NR perturbations.
 7 Conclusion
-In this paper, we evaluate the extent of GPT-4’s verbatim memory for fictional works and its impact on downstream tasks.
-We also propose methods to mitigate this impact and encourage the model to use reasoning.
-We draw several key conclusions: (1) The model always tends to first rely on verbatim memory when performing character understanding tasks, and it consistently benefits from this approach.
-(2) When the predicted text differs significantly from the pretraining data, the model is more likely to use reasoning to attempt to compensate.
+In this work, we introduce the concept of gist memory and design prompts for six types of character understanding tasks, enabling models to reduce reliance on memorization and enhance reasoning-based character interpretation.
+Our approach provides a systematic framework to evaluate the extent to which different character understanding tasks depend on memorization, offering a more accurate assessment of model performance by disentangling genuine reasoning from memorization-driven responses.
+These findings highlight the limitations of current benchmarks and provide insights for designing more robust character understanding evaluations.
 8 Limitations
@@
 Table 5: An example for source prediction task.
-Appendix C Self-generated Character Description
-For the self-generated character description, we use following prompts to ask ChatGPT to generate a brief character summarization for each character.
-The following block-quoted text shows the output we got; importantly, it demonstrates the level of deep knowledge about each of these characters that is baked into the model itself.
-Input: Give me a detailed character description for all main characters in the TV show Friends.
-Output: Rachel Green (Jennifer Aniston):
-Background: Initially introduced as a runaway bride, Rachel starts as a waitress at Central Perk and evolves into a successful fashion executive.
-Character Traits: Fashion-conscious, independent, initially somewhat spoiled, but grows into a more self-sufficient and career-focused woman.
-Romantic Arc: Has an iconic on-again, off-again relationship with Ross throughout the series.
-Ross Geller (David Schwimmer):
-Profession: Paleontologist working at a museum.
-Character Traits: Intelligent, nerdy, and somewhat socially awkward. Ross has a complicated love life and is often portrayed as the "smart" but emotionally vulnerable member of the group.
-Romantic Arc: His tumultuous relationships, particularly with Rachel, are central to the series.
-Monica Geller (Courteney Cox):
-Profession: Chef and later a restaurant owner.
-Character Traits: Competitive, organized, and slightly obsessive about cleanliness. Monica is Ross’s sister and becomes the motherly figure of the group.
-Romantic Arc: Her relationship with Chandler Bing becomes a major storyline.
-Chandler Bing (Matthew Perry):
-Profession: Holds an office job with a mysterious title related to statistical analysis.
-Character Traits: Known for his sarcastic humor, commitment issues, and use of humor as a defense mechanism. Evolves into a more mature and emotionally secure character.
-Romantic Arc: Develops a significant romantic relationship with Monica.
-Joey Tribbiani (Matt LeBlanc):
-Profession: Aspiring actor throughout the series.
-Character Traits: Easygoing, simple-minded, and often portrayed as the least academically inclined. Known for his catchphrase "How you doin’?" and his love for food.
-Romantic Arc: Has numerous short-term relationships, and his genuine friendship with Chandler is a key dynamic.
-Phoebe Buffay (Lisa Kudrow):
-Profession: Eccentric musician with a unique and mysterious past.
-Character Traits: Quirky, free-spirited, and spiritually inclined. Phoebe often provides offbeat humor and memorable songs.
-Romantic Arc: Her romantic life includes unique relationships, and her eccentricity adds a distinctive element to the group.
-These six characters form the core group of friends in the series, and their individual personalities, relationships, and growth over the ten seasons contribute to the show’s humor, heart, and enduring popularity.
+Appendix C Extended Dataset Analysis
+For the six representative character understanding tasks studied in our work, we selected one dataset for each category.
+However, we also recognize that there are many different benchmark datasets for each type of task.
+When selecting representative datasets for each category, we focused on choosing those with more diverse sources.
+Additionally, we supplemented the selection with eight datasets of the same type but derived from different sources in Table 6.
+Task Name Task Type Source
+Character Mining Character Identification TV Shows
+LISCU Character Identification Book Reviews
+MovieQA Visual Question Answering Movies
+Liter Coreference Character Coreference Novels
+HVV Detect Role Detection Movies, News
+SYMON Visual Question Answering Movies
+CEB Character Representation Novels
+Love and Violence Summarization TV Shows
+Table 6: Overview of 14 tasks in the character understanding benchmark.
 Appendix D Prompts for GPT4
-for personality traits extraction ’you are a psychology research assistant designed to help analysis character personalities according to conversations.’ conversation ’Here ends the conversation.Give me a character description for each main character according to this conversation.’
+for personality traits extraction ’you are a psychology research assistant designed to help analysis character personalities according to conversations.’ conversation ’Here ends the conversation.Give me a character description for each main character according to this conversation.’ Character Description Here is a character description:Monica: Responsible, caring, and organized.
+Assertive and confident in her actions.Chandler: Witty and self-deprecating with an approachable sense of humor.
+Exhibits insecurity and anxiety in his dialogues, making references to uncomfortable situations and questioning his own actions.Rachel: Spontaneous and open to change, she takes risks and is adaptable.
+She is also reliant on her relationships with others, signifying her dependency and need for support.’ ’Joey: Energetic, extroverted, and casual.
+Lacks the sensitivity of others feelings at times but genuinely care about friends.Phoebe: Quirky, eccentric, and a free spirit.
+Her train of thought tends to lean toward the unusual and bizarre.
+However, she is also compassionate and caring.Ross: Insecure and somewhat neurotic and vulnerable.
+His behavior is indicative of someone who is going through emotional turmoil.’
 Appendix E Name Replacement and Character Comparison
@@
 The six selected benchmarks cover a range of character-centric tasks, designed to test different aspects of character understanding(Table 1):
-  • TVShow Guessing: Based on the TVSHOWGUESS dataset, which contains scripts from five popular TV series: Friends, The Big Bang Theory, The Office, Frasier, and Gilmore Girls Sang et al.
+  • TVShow Character Guessing: Based on the TVSHOWGUESS dataset, which contains scripts from five popular TV series: Friends, The Big Bang Theory, The Office, Frasier, and Gilmore Girls Sang et al.
 (2022c), this is a multi-choice classification task.
@@
 The model is required to generate a coherent summary from dialogue utterances and scene descriptions, capturing key plot points while excluding irrelevant details.
+Appendix H Gist Prompting Strategies Across Tasks
+We design Gist Prompting methods to reduce reliance on memorization and promote inference-based character understanding. Below, we detail the task-specific prompt formulations.
+H.1 Task-Specific Gist Prompts
+Character Guessing
+"Analyze the dialogue and infer the most likely speaker based on their speech patterns, personality traits, and relationships with other characters.
+Do not rely on direct recall of specific names but rather use reasoning from the provided context."
+Coreference Resolution
+"Identify which mentions in the text refer to the same character by analyzing linguistic patterns, contextual cues, and discourse coherence.
+Do not rely on pre-learned entity associations but rather resolve coreference based on narrative structure."
+Personality Understanding
+"Based on the character’s dialogue and actions, infer their personality traits.
+Focus on how they interact with others, their tone, and behavioral tendencies.
+Avoid using prior memorized descriptions of characters and derive insights purely from the given passage."
+Role Detection
+"Determine whether a character plays a key role in the narrative (e.g., antagonist, victim, or protagonist) by analyzing their function within the story.
+Identify how their actions and relationships define their role rather than relying on predefined labels."
+Open-Domain QA
+"Answer the question using only the given passage.
+Identify relevant information by reasoning over the text and analyzing interactions between characters.
+Do not rely on external knowledge or prior memorized facts."
+Summarization
+"Summarize the passage by identifying key events, character interactions, and narrative progression.
+Construct a coherent summary that captures the essence of the scene without directly copying text segments."

### v5

@@
 Abstract
-Recently, Large Language Models (LLMs) have shown impressive performance in various character understanding tasks.
-However, the extensive pre-training corpora used by LLMs raise concerns that they may rely on memorizing popular fictional works rather than genuinely understanding and reasoning about them.
-Prior work has addressed memorization in LLM reasoning tasks, but these approaches do not directly apply to character understanding.
-To address this, we leverage the cognitive science concepts of "gist memory" and "verbatim memory", where verbatim memory corresponds to exact recall, while gist memory captures underlying meaning.
-We design a hard-soft intervention framework, where the hard setting disrupts memorization by perturbing key entities, while the soft setting maintains the original text but guides the model toward reasoning.
-Our experiments across six character understanding tasks with GPT-4o, LLaMA3.3-70B, and DeepSeek V3 demonstrate that reducing verbatim memory reliance shifts model behavior toward reasoning.
-Our method leads to a performance drop of up to 45.8% when models can no longer depend on memorization, highlighting the extent to which existing benchmarks inflate performance through data contamination.
-Additionally, we show that prompting techniques can systematically alter a model’s response strategy, making it more dependent on either memorization or reasoning.
+Large Language Models (LLMs) achieve striking performance on character understanding tasks, yet this success raises a key concern: are they truly reasoning about fictional characters, or merely recalling them from memorization?
+While prior work has explored memorization in general reasoning tasks, it remains unclear how memorization affects character understanding.
+To investigate this, we draw on cognitive science distinctions between "verbatim" (exact recall) and "gist" (semantic abstraction) memorization, and introduce a graded intervention framework with two levels of disruption.
+The soft setting preserves surface text but applies prompting strategies to encourage reasoning over recall, while the hard setting perturbs key character references to directly block memorization cues.
+Evaluating GPT-4o, LLaMA3.3-70B, and DeepSeek V3 on six character understanding benchmarks, we find that our hard setting leads to a dramatic performance drop of 45.2%—revealing substantial reliance on verbatim memorization and data leakage in existing benchmarks.
+In contrast, the soft setting steers models to reason more robustly without harming performance, showing that prompting alone can shift LLMs between memorization-based and reasoning-based strategies.
 Memorization Over Reasoning? Exposing and Mitigating Verbatim Memorization in Large Language Models’ Character Understanding Evaluation
@@
 1 Introduction
-Character understanding requires the models to grasp the characters’ personalities and traits while reading a story, which aims to evaluate the models’ capability to comprehend personalities, goals, roles, and relationships of various fictional characters Yu et al.
+Character understanding tasks assess a model’s ability to comprehend fictional characters’ personalities, goals, roles, and relationships within a narrative Yu et al.
 (2023a); Vallurupalli et al.
@@
 (2021).
-However, the impressive performance of LLMs on such benchmarks Li et al.
+Recent LLMs have shown strong performance on such benchmarks Li et al.
 (2023); Stammbach et al.
 (2022); Goyal et al.
-(2023) raises the question: does it reflect genuine understanding or merely strong memorization Chang et al.
+(2023), raising an important question: do these models truly understand characters, or are they simply recalling text seen during pretraining Chang et al.
 (2023a)?
-For example, in the TV show speaker guessing task shown in Figure  1, the model is expected to identify Sheldon based on its memory of the character from previous episodes (e.g., preparing lessons for Howard).
-However, if the model simply memorizes the exact words, it could make predictions based on a very short sentence by directly extracting the speaker information from it.
-To truly stimulate a model’s character understanding ability, it is crucial to understand how the model may be generating its responses.
-This can be especially challenging given the extensive and opaque pre-training corpora used by LLMs, making it particularly severe for character understanding tasks, as many benchmarks are derived from well-known and widely consumed fiction or scripts Sang et al.
+For instance, in a speaker guessing task from The Big Bang Theory (Figure 1), a model may identify Sheldon correctly not by reasoning over past actions (e.g., preparing lessons for Howard), but by matching exact phrases seen in previous data.
+This concern is especially relevant in character understanding, as many benchmarks are based on widely consumed fictional works, such as TV shows, movies, or novels Sang et al.
 (2022a).
-We hypothesize that this can lead to a form of memorization called “verbatim memory,” which encodes precise details and remembers every word exactly Huang et al.
+Given the massive and opaque pretraining corpora of LLMs, models are likely to have encountered these works verbatim.
+As illustrated in Figure 1, such memorization can yield high performance even without genuine understanding.
+As a result, performance may reflect reliance on what cognitive psychology calls verbatim memorization—exact recall of surface text Huang et al.
 (2024); Ippolito et al.
-(2022)—rather than reasoning over recalled overarching patterns (what is referred to as “gist memory” Reyna and Brainerd (1998)).
-Figure 1: This figure illustrates the distinct mechanisms of gist and verbatim memory processes in the speaker-guessing task for a script.
-Verbatim memory (bottom right) aims to recall as much specific information as possible, including who said each sentence.
-If the memorization is not precise enough, the model fails to provide the correct answer.
-In contrast, gist memory (top right) focuses on extracting core information, such as character relationships, personalities, and key events.
-By relying on gist memory, the model is encouraged to engage in reasoning rather than purely relying on memorization.
-Motivated by the concept of gist memory, we propose a prompt-based approach to mitigate model reliance on verbatim memorization in character understanding tasks.
-We evaluate our method on GPT-4o, LLaMA3-70B, and DeepSeek V3, introducing a hard-soft setting framework: the hard setting blocks memorization by replacing character names, while the soft setting retains the original text but employs gist-based prompts to encourage reasoning over recall.
-Our ablation study confirms the effectiveness of both interventions.
-Building on these findings, we extend our approach to five character understanding tasks, designing task-specific gist prompts and evaluating models under both settings.
-Our results indicate that the Gist+NR approach best reduces memorization dependency.
-In summary, we outline our study and key contributions as follows:
+(2022)—rather than gist memorization, which captures semantic understanding and general patterns Reyna and Brainerd (1998).
+This challenges the validity of current evaluations: are models demonstrating true reasoning, or simply relying on verbatim memorization to regurgitate answers?
+In this work, we formulate a testable hypothesis: if LLMs’ strong performance is largely driven by memorized surface cues, then removing those cues—such as character names—should lead to a significant drop in accuracy.
+Figure 1: Gist vs.
+verbatim memorization in speaker identification under original and anonymized scripts.
+This figure compares two memorizaiton strategies in identifying speakers (P0, P1) from a dialogue.
+On the left, character names are intact; on the right, all names are anonymized.
+The gist-based approach (blue) is prompted to "analyze personality traits, tones, and relationships", and infers speaker identity through reasoning—remaining robust even when names are masked.
+In contrast, the verbatim-based approach (yellow) is prompted to "recall names from exact dialogue memorization", and may misattribute the script (e.g., predicting Friends instead of The Big Bang Theory) when superficial cues are removed.
+To investigate this hypothesis, we develop a graded intervention framework with two levels of disruption: a soft setting and a hard setting.
+The soft setting retains the original input scripts but uses gist-based prompting to guide the model toward abstraction and inference, without removing surface-level cues.
+In contrast, the hard setting replaces all character names with culturally distinct alternatives, aiming to block verbatim memorization while preserving the narrative content.
+We apply this graded intervention framework to six character understanding tasks and evaluate three LLMs: GPT-4o, LLaMA3-70B, and DeepSeek V3.
+Our experiments provide clear support for our initial hypothesis: removing verbatim memorization cues significantly reduces LLMs’ performance on character understanding tasks, confirming that models heavily rely on memorized surface content.
+Specifically, we observe performance drops of up to 45.8% when verbatim memorization cues are blocked.
+In contrast, gist-based prompting maintains accuracy by steering model behavior toward reasoning rather than verbatim recall.
+These results suggest that current benchmarks may overestimate models’ genuine reasoning capabilities and that prompting strategies can effectively disentangle reasoning from memorization.
+In summary, our key contributions are:
   1. 1.
-    Character understanding tasks heavily rely on verbatim memorization.
-Across six benchmark tasks, all evaluated LLMs show significant performance drops when explicit memorization is disrupted, confirming that these tasks often assess recall rather than true reasoning (Table 6.2).
+    Identifying Memorization in Character Understanding: We empirically demonstrate significant performance drops (up to 46.2%) across benchmarks when verbatim memorization cues are disrupted, revealing a heavy reliance on surface-level recall rather than genuine reasoning.
   2. 2.
-    Gist-based prompting effectively reduces memorization dependency.
-Introducing gist-based prompts systematically lowers reliance on verbatim recall across all tasks, encouraging models to use reasoning instead of mechanical memorization (Section 6).
+    A Graded Framework to Disentangle Reasoning from Memorization: Inspired by cognitive distinctions, we introduce a soft–hard intervention framework.
+The soft setting employs gist-based prompting to encourage abstraction, while the hard setting explicitly blocks verbatim recall through surface-level perturbation (Section 6).
   3. 3.
-    Verbatim and gist-based prompting modulate model response strategies.
-Models adjust their response strategies based on the prompting style—verbatim-based prompts reinforce reliance on memorization, while gist-based prompts encourage reasoning-driven answers (Section 4.3).
+    Prompting as a Diagnostic Tool: We show that prompting style profoundly influences model behavior: verbatim prompts reinforce memorization, whereas gist prompts facilitate deeper inference, providing a practical approach to probe underlying reasoning processes in LLMs (Section 4.3).
 2 Related Work
-2.1 Character Understanding and its evolution with LLM
-Character understanding has made significant contributions to advancing story comprehension in psychological, literary, and educational research  Bower and Morrow (1990); McKee (2010); Paris and Paris (2003).
-This multifaceted field encompasses a variety of tasks, including character recognition, coreference resolution, summarization, role detection, goal understanding, personality analysis, and question answering Sang et al.
-(2022b); Bamman et al.
+2.1 Character Understanding and Its Evaluation with LLMs
+Character understanding has long been a focus in psychological, literary, and educational research Bower and Morrow (1990); McKee (2010); Paris and Paris (2003).
+It spans tasks such as character recognition, coreference resolution, summarization, role detection, goal modeling, personality analysis, and question answering Bamman et al.
 (2020); Chen et al.
@@
 (2023b).
-Recent studies have shown promising performance by LLMs on character understanding tasks Li et al.
+Recent work demonstrates that LLMs achieve strong results on character-centric benchmarks Li et al.
 (2023); Stammbach et al.
@@
 (2023).
-However, these studies often overlook the potential influence of memorization on model performance.
-Fictional works, including movie and TV show scripts, famous novels, and news texts, serve as primary sources for narrative understanding Sang et al.
+However, many of these datasets are derived from publicly available or widely consumed fictional sources, such as movie scripts, TV shows, and novels Sang et al.
 (2022a); Tapaswi et al.
-(2016); Caselli and Vossen (2017).
-These texts, much like copyrighted books Chang et al.
-(2023a), are prone to strong memorization.
-Naturally, this raises curiosity about how memorization affects downstream tasks, particularly character understanding.
-While previous research has demonstrated the memorization capabilities of LLMs, our work further explores its specific impact on character understanding and proposes an approach to mitigate this effect.
-2.2 LLMs’ Memorization and Reasoning
-Among LLMs’ various capabilities Li et al.
-(2024c, e); Zhang et al.
+(2016); Caselli and Vossen (2017), which makes them vulnerable to pretraining overlap and surface-level memorization Chang et al.
+(2023a).
+Although prior research has studied LLM memorization broadly Carlini et al.
+(2023); Xie et al.
+(2024); Yang et al.
+(2023a), little is known about how it specifically affects character understanding tasks.
+Our work fills this gap by examining memorization’s role in these tasks and proposing interventions to distinguish between recall and reasoning.
+2.2 Memorization and Reasoning in LLMs
+LLMs’ impressive reasoning capabilities have attracted significant attention Tong et al.
+(2023); Yang et al.
+(2023b); Xiong et al.
+(2024); Tan et al.
+(2024); Jiang et al.
+(2025), yet their behavior is also strongly influenced by memorization Zhou et al.
+(2023); Chen et al.
+(2024); Biderman et al.
+(2024).
+Prior studies indicate that memorization of training data can artificially inflate benchmark results Khandelwal et al.
+(2020); Blevins and Zettlemoyer (2022), while conflicting memorized content can negatively affect performance Su et al.
+(2024); Wang et al.
+(2024); Yang et al.
+(2025).
+Unlike structured reasoning tasks, character understanding tasks inherently involve some memorization (e.g., narrative consistency), complicating the distinction between genuine reasoning and mere recall.
+Cognitive science provides a useful conceptual framework for analyzing memory, distinguishing between verbatim memorization—the precise recall of surface-level details—and gist memorization—the abstraction of semantic meaning Reyna and Brainerd (1998); Brainerd and Reyna (2002).
+To avoid confusion with architectural notions of "memory" in NLP and machine learning Alizadeh et al.
+(2024); Kwon et al.
+(2023), we consistently adopt these terms to describe different recall styles rather than neural memory mechanisms.
+In our work, this distinction serves as the theoretical foundation for disentangling shallow memorization from deeper reasoning behavior in large language models.
+Recent NLP research has formalized methods for quantifying and disentangling memorization from reasoning Lesci et al.
+(2024); Jin et al.
+(2024); Carlini et al.
+(2023); Prabhakar et al.
+(2024), examining phenomena such as multifaceted memorization Prashanth et al.
 (2024); Li et al.
-(2024d), the reasoning abilities Tong et al.
-(2023); Yang et al.
-(2023); Xiong et al.
-(2024b); Tong et al.
-(2024); Xiong et al.
-(2024a) of LLMs Li et al.
-(2024b); Tan et al.
-(2024) have garnered significant attention, but their memorization has also been shown to play a crucial role in influencing benchmark reasoning performance Zhou et al.
-(2023); Xie et al.
-(2024a); Chen et al.
-(2024); Li et al.
-(2024a); Biderman et al.
+(2024) and logit-based tracing Chen et al.
 (2024).
-The influence of memorization is complex: data contamination can inflate performance metrics Khandelwal et al.
-(2020); Blevins and Zettlemoyer (2022); Carlini et al.
-(2023), while knowledge conflict can reduce them Su et al.
-(2024); Wang et al.
-(2024); Xie et al.
-(2024b); Xu et al.
-(2024).
-While prior work has focused on detecting and mitigating memorization effects in reasoning tasks such as numerical, commonsense, and logical reasoning Xie et al.
-(2024a); Sainz et al.
-(2023); Li et al.
-(2024f), character understanding tasks introduce additional challenges due to their diverse formats and varying memory demands.
-Unlike structured reasoning tasks where excessive memorization is clearly undesirable, character understanding often requires a nuanced balance between recall and inference.
-2.3 Verbatim and Gist Memory in LLM Reasoning
-Cognitive science distinguishes between verbatim memory, which retains exact details, and gist memory, which captures the essential meaning Brainerd and Reyna (2002); Reyna and Brainerd (1998).
-For example, verbatim memory would store a sentence like "The black-and-white cat sat on a soft, red mat in the living room" with all details intact, while gist memory would reduce it to "A cat was sitting indoors," focusing on the core idea while discarding specifics.
-Human reasoning primarily relies on gist memory, as the two types of memory function independently rather than hierarchically Brainerd and Reyna (1992).
-In LLM research, memorization is often equated with verbatim memory, where models retrieve and reproduce exact strings from pretraining data Carlini et al.
-(2023).
-A common mitigation strategy involves curating test samples that the model has never seen before Xie et al.
-(2024a).
-However, in character understanding tasks, acquiring obscure or newly released datasets can be costly and difficult to control in terms of quality.
-Instead, our work proposes leveraging gist memory to reduce reliance on precise memorization, encouraging LLMs to generalize character understanding beyond string-matching.
-By shifting the focus toward gist memory, we aim to align LLM character analysis with human-like reasoning rather than dataset recall.
+Mitigation strategies (e.g., out-of-distribution filtering Biderman et al.
+(2024)) are challenging to apply to character-centric tasks, which typically rely on known fictional content.
+Thus, we propose a prompting-based approach designed specifically to reduce verbatim reliance and promote gist-level reasoning in character understanding tasks.
 3 Data and Experiment Setup
-In our work, we utilize two distinct datasets to evaluate LLMs on: (1) their reasoning capabilities in character understanding under gist memory and (2) their memorization of original source texts.
-3.1 Benchmark Datasets for Character Understanding Tasks
-To evaluate the interplay between memorization and reasoning in character understanding tasks, we employ six publicly available datasets covering diverse task types.
-These datasets span different narrative mediums, including TV shows, movies, and novels, ensuring diversity in both content and character representation.
-The task formulations and dataset compositions are detailed below, with an overview provided in Table 1.
-Further task details are available in  6.1.
-Datasets
-We evaluate character understanding using six publicly available datasets spanning diverse sources, including TV shows, movies, and novels.
-For coreference resolution, we use the latest release of the MovieCoref dataset  [^1] https://github.com/usc-sail/mica-character-coref, which provides annotated coreference chains in movie scripts.
-For character guess, we adopt TVSHOWGUESS  [^2] https://github.com/YisiSang/TVSHOWGUESS, where models must identify the speaker of a given dialogue.
-For personality understanding, we use PERSONET  [^3] https://github.com/Gorov/personet_acl23, a dataset derived from literary sources with personality trait annotations.
-For role detection, we employ CSI  [^4] https://github.com/EdinburghNLP/csi-corpus, which focuses on identifying key characters in crime narratives.
-For open-domain question answering, we evaluate models on FriendsQA  [^5] https://github.com/emorynlp/FriendsQA, where questions are answered by extracting spans from TV show dialogues.
-Lastly, for summarization, we use SummScreen  [^6] https://github.com/mingdachen/SummScreen, which pairs TV show transcripts with corresponding summaries.
-We follow the original dataset authors’ preprocessing, development, and test splits to ensure comparability.
-Further dataset details are provided in Appendix G.
+To comprehensively evaluate character understanding in LLMs, we compile two datasets: CharBench and CharScript.
+Inspired by NLP benchmark collections such as GLUE Wang et al.
+(2018), our CharBench integrates six publicly available datasets spanning character-centric tasks like coreference resolution, personality inference, and dialogue-based QA (details in Section 6.1).
+The second dataset, CharScript, is newly constructed for this study by systematically sampling textual segments from original fictional sources underlying CharBench, explicitly probing models’ verbatim memorization (details in Section 3.2).
+3.1 CharBench: A Benchmark Collection for Character Understanding
+CharBench aggregates six widely recognized NLP benchmarks covering diverse narrative mediums—TV shows, movies, and novels.
+To ensure comparability, we retain each dataset’s original splits, preprocessing, and evaluation metrics.
+Detailed task descriptions and implementations are provided in Appendix E.
+Benchmark Sources
+Following established NLP benchmark practices Wang et al. (2018), CharBench includes:
+  • MovieCoref Chen and Choi (2016): Coreference resolution in movie scripts.
+  • TVSHOWGUESS Sang et al. (2022b): Speaker identification in TV dialogues.
+  • PERSONET Yu et al. (2023b): Personality inference from literary passages.
+  • CSI Corpus Frermann et al. (2018): Role detection from crime shows.
+  • FriendsQA Yang and Choi (2019): Multi-turn dialogue-based QA.
+  • SummScreen Chen et al. (2022): TV episode summarization.
 Evaluation Metrics
-We followed the original evaluation metrics used in the respective tasks, shown in Table 1.
-Task Name Task Type Data Source Metrics
+We strictly follow original dataset metrics (exact match, ROUGE, F1 score), as detailed in Table 1.
+Task Name Task Type Data Source Metric
 TVShow Guess Character Guessing TV Shows Exact Match
@@
 Table 1: Overview of benchmark tasks, including their type, data sources, and evaluation metrics.
-3.2 Fictional Works Memorization Benchmark Dataset
-This dataset is designed to evaluate LLMs’ memorization of fictional scripts by assessing their ability to recall verbatim details from diverse narrative sources.
-Benchmark Sources
-The six benchmark datasets described in Section 6.1 are derived from a diverse set of sources, including 91 TV shows, 12 movies, and 33 classic literary works.
-Dataset Construction
-To ensure broad coverage across different narrative mediums, we sample 200 script segments from 40 selected sources, comprising 11 TV series, 9 movies, and 20 novels.
-Each source contributes 5 excerpts, with an average segment length of 312 words.
-Segment Format
-Each excerpt follows the format: TV Show Title; Scene, where the scene consists of a mixture of dialogue and narration extracted from the original script.
+3.2 CharScript: Fictional Works Memorization Benchmark Dataset
+Motivated by recent NLP research on memorization in language models Chang et al.
+(2023b), we constructed the CharScript dataset to evaluate whether LLMs can recognize the source of fictional text segments.
+Following a similar setup to Chang et al.
+(2023b), we probe whether models can correctly attribute short passages to their original narrative sources, as an indicator of verbatim memorization.
+We present our experimental design and findings in Section 5.1.
+We systematically sampled and extracted 200 textual segments, averaging approximately 312 words each, from 40 diverse fictional sources (11 TV series, 9 movies, and 20 novels).
+These sources were carefully selected to maximize narrative diversity, textual complexity, and representativeness of popular fictional works included in CharBench.
+Each source contributes exactly 5 segments, structured uniformly in the format [Title]; Scene context, combining both original dialogue and narrative descriptions.
 3.3 Models
-we evaluate both open-source and closed-source models, leveraging their APIs for inference.
-Our evaluation includes three frontier LLMs: the open-source Llama 3.3-70B AI (2024) and DeepSeek V3 Liu et al.
-(2024), as well as the closed-source GPT-4o OpenAI (2023).
+We evaluate three frontier LLMs—open-source LLaMA 3.3-70B AI (2024) and DeepSeek V3 Liu et al.
+(2024), as well as the closed-source GPT-4o OpenAI (2023)—via API-based inference.
+To address stochastic variability, each model was run three times per setting, and average performance is reported.
+The total cost of our experiments was approximately \$100.
 4 Soft Setting: Memorization-Controlled Prompting in Character Guessing
-First, we investigate whether prompting strategies can subtly influence LLMs’ response patterns in a soft setting, where the model is not explicitly restricted from using memorization but is instead encouraged to rely more on reasoning over direct retrieval.
+In the introduction, we formulated a testable hypothesis that LLMs’ strong performance on character understanding tasks relies significantly on verbatim memorization.
+As a first step toward testing this hypothesis, we investigate whether prompt design alone—without explicitly removing surface-level textual cues—can shift model responses from verbatim recall toward gist-based reasoning.
 4.1 Character Guessing Task Format
-Let a dialogue segment be represented as $D=\{U_{1},U_{2},...,U_{n}\}$, where each utterance $U_{i}$ is paired with a speaker label $S_{i}$.
-In the Character Guessing Task, the model is given a dialogue excerpt without speaker annotations and must predict the correct speaker for each utterance, $f_{\text{guess}}(U_{i})\rightarrow S_{i}$.
-We follow the original format and setting Sang et al.
-(2022b), using the TVShow Guess Dataset described in Section 3.1.
-This task is designed to prompt the model to make guesses based on character relationships, key events, and personality traits within the text.
-However, its format closely resembles the name cloze tests Chang et al.
-(2023b) recently used to evaluate LLM memorization.
+We select the Character Guessing Task specifically because it closely mirrors the name cloze tests Chang et al.
+(2023b) previously established as effective probes for verbatim memorization in LLMs.
+Name cloze tests measure memorization by removing character names from familiar contexts and testing if models can accurately recall them.
+Similarly, in our task, speaker labels are removed from dialogue excerpts, requiring the model to infer speakers solely from textual context.
+If models rely heavily on verbatim memorization, accuracy should significantly degrade when surface cues—such as explicit character names—are disrupted.
+Conversely, if models genuinely reason about character interactions and relationships, performance should remain relatively stable despite these perturbations.
+Following the original TVShow Guess setup Sang et al. (2022c), the model is tasked with predicting the speaker of each utterance in a dialogue segment where all speaker labels are removed.
 4.2 Prompting Strategies
-We design two distinct prompting conditions to examine how LLMs rely on different memory mechanisms when identifying characters.
+We design two distinct prompting conditions to examine how LLMs rely on different memorization mechanisms when identifying characters.
 Verbatim prompting encourages mechanical memorization, where the model recalls exact substrings from its training data and reconstructs missing parts without reasoning.
-In contrast, gist prompting focuses on extracting key information and performing memory-based matching.
+In contrast, gist prompting focuses on extracting key information and performing memorization-based matching.
 For character understanding tasks, the most relevant gist cues typically include character relationships, significant events, and personality traits.
@@
 The two variations are as follows:
-Verbatim Prompt: "…Do this by direct memory retrieval—do not infer based on traits, interactions, or context. Recall names exactly from past dialogue knowledge without explanations or reasoning."
-Gist Prompt: "…Do this by analyzing relationships, key events, and personality traits to match them with known characters. Do not directly retrieve from memory who said the sentences."
+Verbatim Prompt: "…Do this by direct memorization retrieval—do not infer based on traits, interactions, or context.
+Recall names exactly from past dialogue knowledge without explanations or reasoning."
+Gist Prompt: "…Do this by analyzing relationships, key events, and personality traits to match them with known characters. Do not directly retrieve from memorization who said the sentences."
 4.3 Result and Analysis
-Figure 2 visualizes the performance of the three models on the Character Guessing task under different prompting conditions.
-The performance trends remain consistent across models, and our key observations are as follows:
-  • Baseline Performance: The baseline condition reflects the overall performance of the models, with GPT-4o demonstrating a clear lead, while Llama 3.3-70B and DeepSeek V3 exhibit similar performance.
+Our key observations are as follows:
+  • Baseline Performance: The baseline condition reflects the overall performance of the models, with GPT-4o demonstrating a clear lead to others, while Llama 3.3-70B and DeepSeek V3 exhibit quite similar performance.
   • Memorization Reliance: All three models show that the Verbatim condition performs nearly identically to the baseline, supporting our hypothesis that models primarily rely on memorization when answering the Character Guessing task.
-  • Effect of Gist Prompting: The Gist condition results in a non-trivial performance drop, suggesting that our prompt effectively reduces the models’ reliance on memorization when generating responses.
+  • Effect of Gist Prompting: The Gist condition results in a non-trivial performance drop, approximately 10%, suggesting that our prompt effectively reduces the model’s dependence on memorization when generating responses.
+Our results indicate that gist-based prompting can indeed shift model responses toward reasoning, as evidenced by the performance drop under the Gist condition (see Figure 2).
+However, the moderate performance drop (approximately 10%) suggests that even gist prompting may not completely eliminate the models’ reliance on verbatim memorization.
+Previous research similarly indicates that prompting without explicit reasoning steps may fail to fully engage true reasoning processes Yu (2025).
+Therefore, to conclusively test our hypothesis—that models’ strong benchmark performance relies significantly on verbatim memorization—we introduce a more direct and explicit intervention aimed at disrupting memorized surface-level cues.
 Figure 2: The bar chart compares the performance of three LLMs under three prompting settings.
@@
 5 Hard Setting: Memorization-Controlled Prompting under Name Replacement
-Furthermore, we aim to validate whether our prompting strategies can genuinely alter the model’s reasoning patterns under a hard setting, where memorization is explicitly restricted.
-Previous research has suggested that name replacement may impact model performance Yu et al.
-(2022).
-In this section, we confirm this hypothesis by demonstrating that name replacement significantly affects the model’s memorization ability and quantifying its impact.
-Then, we examine how prompting strategies influence LLMs’ reliance on memorization versus reasoning when memorization is constrained by name replacement.
+To effectively suppress verbatim memorization and induce gist-style reasoning, we design a stronger intervention: replacing character names to explicitly remove memorized surface cues.
+While prior work uses name replacement to improve model robustness Yu et al.
+(2022), we take this idea further by systematically testing whether such name-based interventions are strong enough to disrupt verbatim memorization and induce gist-based reasoning.
 5.1 Name Replacement Strategies and Their Impact on Memorization
@@
 To systematically analyze the impact of name replacement on model performance, we introduce three distinct levels of name modifications:
-  • Name Masking: Replacing all character names with placeholders (e.g., Sheldon → P0, Amy → P1).
-  • Cross-Cultural Name Replacement: Substituting names with those from a different cultural background while maintaining gender consistency (e.g., SheldonJie Zhang, AmyXiaoling Wang).
+  • Name Masking: Replacing character names with placeholders (e.g., Sheldon, Amy → P0, P1).
+  • Cross-Cultural Name Replacement: Substituting names with those from a different cultural background (e.g.English → Chinese) while maintaining gender consistency (e.g., Sheldon, AmyJie Zhang, Xiaoling WangSun et al.
+(2024).
 This tests whether models rely on specific name distributions for character inference.
-  • Same-Cultural Name Replacement: Replacing names with other culturally consistent alternatives (e.g., SheldonAndrew, AmySally).
-This retains cultural familiarity while disrupting direct memorization.
+  • Same-Cultural Name Replacement: Replacing names with other culturally consistent alternatives (e.g., Sheldon, AmyAndrew, Sally).
+This retains cultural familiarity while disrupting direct memorization by name replacements.
 Memorization Evaluation Task Format
-We design the Source Identification Task where the model is given a dialogue excerpt without source annotations and must predict the correct origin of the text.
-Let a dialogue segment be represented as $D=\{U_{1},U_{2},...,U_{n}\}$, where each utterance $U_{i}$ is part of a larger script from a specific source.
-The task requires the model to determine the correct source, $f_{\text{source}}(D)\rightarrow S$, where $S$ denotes the source work.
-We use the Fictional Works Memorization Benchmark Dataset described in Section 3.2.
+To directly quantify how much verbatim memorization models rely on, we design a Source Identification Task.
+This task closely resembles established memorization evaluation methods, such as source attribution or textual cloze tests Chang et al.
+(2023b).
+Specifically, given a dialogue excerpt $D=\{U_{1},U_{2},...,U_{n}\}$ without source annotations, the model must identify the correct origin of the text, predicting $f_{\text{source}}(D)\rightarrow S$, where $S$ is the specific fictional source.
+The rationale behind this design is straightforward: if a model has memorized verbatim text from pretraining, it should accurately identify the source based purely on surface-level textual recall.
+Conversely, if name replacement or prompting disrupts memorization effectively, we expect a significant decrease in source-identification accuracy.
+Thus, observing substantial performance degradation would strongly support the claim that models’ prior accuracy was driven largely by verbatim memorization rather than reasoning.
+We use the CharScript: Fictional Works Memorization Benchmark Dataset described in Section 3.2 to rigorously evaluate this effect.
 Impact of Name Replacement on Memorization.
@@
 Among them, Cross-Cultural Name Replacement consistently induces the strongest forgetting effect across all models, indicating that altering names to those from a different cultural background disrupts memorized associations more effectively than other replacement methods.
+This suggests that cultural distance may play a crucial role in weakening memorized identity-linking patterns in models.
 Figure 3: The heatmap presents accuracy percentages for source identification across three LLMs under different name replacement conditions.
@@
 Different name replacement strategies have shown a significant impact on the memorization ability of LLMs.
-Building on this, we apply the name replacement strategies introduced in Section 5.1 to evaluate the Character Guessing task using the different prompting strategies discussed in Section 4.2.
+Building on this, we apply the name replacement strategies introduced in Section 5.1 to evaluate the Character Guessing tasks, using the verbatim and gist prompting strategies we have discussed in Section 4.2.
 Under the Name Replacement condition, we provide brief character descriptions generated by GPT-4, aiming to enable the model to complete the Character Guessing task through character matching.
-The descriptions used are provided in the Appendix D.
+The descriptions used are provided in the Appendix C.
 To assess the actual effectiveness of these descriptions, we conduct an ablation study.
@@
 Results and Key Findings
-We present the results in Figure 4 and summarize our main findings as follows:
+We present the core results in Figure 4 and summarize four key findings:
   1. 1.
-    Strong Correlation Between Memorization and Performance. Model performance aligns closely with memorization levels, indicating a strong dependence on memorization for accurate predictions.
+    Performance Declines with Reduced Memorization. Accuracy drops as surface cues are removed, revealing the model’s strong reliance on verbatim memorization.
   2. 2.
-    Minimal Impact of Character Descriptions .
-As an ablation study, we evaluate the effect of character descriptions under the Origin Script condition.
-Our results show that incorporating character descriptions does not significantly affect model performance, ruling out potential interference from inaccurate or misleading descriptions.
+    Character Descriptions Have Minimal Impact. Adding character descriptions does not notably change performance, suggesting limited influence on model decisions.
   3. 3.
-    Verbatim Prompting Enhances Memorization.
-Across all three name replacement strategies, Verbatim prompting consistently improves model performance over the baseline, demonstrating that prompting can reinforce memorization and consequently boost accuracy.
+    Prompting Style Matters. Verbatim prompts boost performance by reinforcing memorization, while Gist prompts suppress it and lead to lower accuracy.
   4. 4.
-    Gist Prompting Reduces Reliance on Memorization.
-Performance under the Gist condition declines compared to the baseline, confirming that our Gist prompt effectively reduces the model’s reliance on memorization.
-  5. 5.
-    Name Replacement + Gist Prompting as an Effective Evaluation Strategy.
-We conclude that combining Cross-Culture Name Replacement with Gist prompting provides a reasonable and effective method for evaluating a model’s true capabilities while minimizing reliance on memorization.
+    Name Replacement + Gist Prompting Is Effective. This combination best reduces memorization reliance and reveals the model’s true reasoning ability.
 6 Evaluating Character Understanding Tasks Under Memorization Constraints
-In this section, we evaluate different types of character understanding tasks using the Gist Prompting with Cross Culture Name Replacement approach.
-We first introduce the design principles for Gist Prompts across various task types, followed by an analysis of the results obtained from three different LLMs.
+Sections 4 and 5 provided clear empirical evidence supporting our hypothesis: LLMs’ strong performance on character understanding tasks significantly relies on verbatim memorization.
+To further examine the generalizability of our approach and assess how memorization affects performance across diverse task types, we now extend the combination of Gist Prompting and Cross-Cultural Name Replacement to five additional character understanding benchmarks.
 6.1 Benchmark Tasks Description
-We define the five character understanding tasks being benchmarked and provide an illustrative example.
-Let a dialogue segment be represented as $D=\{U_{1},U_{2},...,U_{n}\}$, where $U_{i}$ is an utterance in the conversation.
-Each utterance $U_{i}$ is associated with a speaker label $S_{i}$, and additional annotations depending on the task.
-  • Coreference Resolution: Cluster character mentions $C=\{c_{1},c_{2},...,c_{n}\}$ into groups referring to the same entity, $f_{\text{coref}}(C)\rightarrow\{\text{clusters of }c_{i}\}$.
-  • Personality Understanding: Classify a character’s personality trait $P$ given dialogue context $D$, $f_{\text{personality}}(D,S_{i})\rightarrow P$.
-  • Role Detection: Determine whether a character plays a key role (e.g., perpetrator in a crime narrative), $f_{\text{role}}(S_{i})\rightarrow\{0,1\}$.
-  • Open-Domain Question Answering: Extract the span $A$ in $D$ that answers a given question $Q$, $f_{\text{QA}}(D,Q)\rightarrow A$.
-  • Summarization: Generate a summary $S$ that captures key events and relationships, $f_{\text{summ}}(D)\rightarrow S$.
+We benchmark five types of character understanding tasks spanning coreference, personality inference, role detection, QA, and summarization.
+Each task evaluates the model’s ability to extract or infer character-related information from dialogue.
 Example Dialogue
@@
   • Preserve Narrative Coherence: In summarization and question answering, focus on maintaining logical flow rather than phrase-matching.
-Detailed gist prompts are included in Appendix H.
+Detailed prompts in Appendix F.
 {NiceTabular}
@@
 6.3 Results and Key Findings
-Our main results and findings are summarized as follows:
+Our evaluation reinforces our conclusions regarding memorization versus reasoning, with three key findings:
   1. 1.
-    Significant Performance Drop with Name Replacement: Across all tasks, we observe a substantial decline in performance when applying the name replacement (NR) method in Table 6.2.
-This result indicates that reducing explicit memorization has a strong impact on model performance, confirming our hypothesis that existing character understanding tasks rely heavily on memorization.
-Moreover, it highlights the potential inflation of benchmark scores under standard evaluation settings.
+    Significant Performance Drop with Reduced Memorization: As shown in Table 6.2, applying the name replacement (NR) method leads to a substantial performance decline across all tasks.
+This confirms our hypothesis that current character understanding benchmarks are heavily reliant on explicit memorization, potentially inflating standard evaluation scores.
+Furthermore, prompting models to use gist reasoning induces an additional performance drop, underscoring the effectiveness of gist prompting in mitigating reliance on surface-level memorization.
   2. 2.
-    Further Degradation with Gist Prompting: The additional performance drop induced by Gist prompting demonstrates its effectiveness across various character understanding tasks.
-This suggests that Gist prompting can further encourage models to rely less on memorization, reinforcing our findings on the role of memorization in these tasks.
+    Performance Gaps across Task Types (Motivation vs.
+Action Prediction): We observe clear differences in memorization effects between action-oriented tasks (e.g., Character Guessing, CSI Role Detection) and motivation-oriented tasks (e.g., Personality Understanding).
+Action prediction tasks typically require models to output discrete labels—such as a character name or role—making it easy for models to memorize associations between surface cues and labels.
+For instance, phrases consistently linked to characters like Sheldon (“I’m not crazy, my mother had me tested”) are easily memorized, but become brittle under name replacement.
+In contrast, motivation prediction tasks involve deeper contextual inference about characters’ internal states and intentions, rendering surface-level memorization less effective, thus promoting genuine reasoning.
   3. 3.
-    Memorization Can Be Detrimental to Character Understanding: Interestingly, we observe performance gains in certain cases under the combined NR+Gist setting.
-Our error case analysis reveals that these improvements stem from resolving knowledge conflicts—instances where the original memorization-based responses misled the model, whereas reducing memorization mitigated such errors, leading to better predictions.
-  4. 4.
-    Consistent Trends Across Models: As shown in Figure 5, despite variations in absolute performance, different models exhibit similar trends in performance degradation across all tasks.
-This consistency suggests that various task designs interact similarly with memorization, reinforcing the broader implications of our findings.
-Figure 5: Performance degradation across LLMs due to the effects of GIST+NR perturbations.
+    Memorization Can Be Detrimental: Interestingly, certain tasks show performance improvements under the NR+Gist setting.
+Our error analysis reveals that reducing memorization resolves knowledge conflicts—cases where memorized but incorrect associations previously misled models, indicating that excessive memorization can actively impair true character understanding.
+These findings are consistent across multiple models (Figure 5), suggesting generalizability and broader implications of memorization biases in current benchmarks.
+Figure 5: Performance degradation across LLMs due to GIST+NR.
+Action prediction tasks (e.g., Guessing, CSI) show greater drops than motivation prediction tasks (e.g., PERSONET, FriendsQA), reflecting their differential reliance on memorization.
 7 Conclusion
-In this work, we introduce the concept of gist memory and design prompts for six types of character understanding tasks, enabling models to reduce reliance on memorization and enhance reasoning-based character interpretation.
-Our approach provides a systematic framework to evaluate the extent to which different character understanding tasks depend on memorization, offering a more accurate assessment of model performance by disentangling genuine reasoning from memorization-driven responses.
-These findings highlight the limitations of current benchmarks and provide insights for designing more robust character understanding evaluations.
+In this work, we introduce the distinction between verbatim memorization and gist memorization, and systematically demonstrate that many existing character understanding evaluations rely heavily on verbatim memorization, potentially inflating perceived reasoning capabilities of LLMs.
+To rigorously test and mitigate this reliance, we propose a novel, graded intervention framework combining Gist Prompting and Cross-Cultural Name Replacement.
+Our experimental results across six diverse character-centric benchmarks clearly show that reducing verbatim memorization cues significantly affects model performance, highlighting the importance of explicitly disentangling memorization from genuine reasoning.
+Our findings emphasize the critical need for developing memorization-aware evaluations, particularly in domains like character understanding where memorization effects are pronounced.
+Broadly, our methodology offers generalizable insights applicable to a wide range of inference tasks where accurately distinguishing between memorization and reasoning is essential.
 8 Limitations
@@
 Our study can be extended by future research in a number of ways.
-First, though we ground and motivate our work through “verbatim” and “gist” memory, we note that any apparent “reasoning” of AI systems may differ fundamentally from those of humans, and we did not pursue human studies to examine this in detail.
+First, though we ground and motivate our work through “verbatim” and “gist” memorization, we note that any apparent “reasoning” of AI systems may differ fundamentally from those of humans, and we did not pursue human studies to examine this in detail.
 Further investigation is needed to explore this topic.
 Second, the relationship between memorization and reasoning remains a subject of debate in cognitive science. Future studies may uncover insights into this relationship from entirely new perspectives.
+Thirdly, While our interventions reveal strong trends, we acknowledge that results may be sensitive to the surface form of prompts.
+Future work could complement our findings with uncertainty-based analyses such as log probability comparisons.
 Finally, our study was centered on select, recent (but not contemporaneous) majority Western popular culture references.
@@
 This work uses the OpenAI API on experiments above, and at no point do we access, or attempt to access, the true training data behind these models, or any underlying components of the systems.
-Risks The several datasets and fictional works scripts in our experiment are sourced from publicly available sources.
-However, we cannot guarantee that they are devoid of socially harmful or toxic language.
+Risks The several datasets and fictional work scripts in our experiment are sourced from publicly available sources.
+However, we cannot guarantee that they are free from socially harmful or toxic language.
 We use ChatGPT  [^1] https://chatgpt.com/ to correct grammatical errors in this paper.
 Appendix A Example of Script-based Character Understanding Task
-In Table 4 we show a full example of the character understanding task. In this example, we have automatically replaced the individual character names with placeholders like P0 and P1.
+In Table˜4 we show a full example of the character understanding task. In this example, we have automatically replaced the individual character names with placeholders like P0 and P1.
 Input P0: Hey, sorry about that P1: No, we’re sorry.
@@
 Table 4: An example from character guessing task.
-Appendix B Example for Verbatim Memory
-B.1 Source Prediction
 Input Tell me the source for following script:Penny: Hey, sorry about that Amy: No, we’re sorry.
@@
 Table 5: An example for source prediction task.
-Appendix C Extended Dataset Analysis
+Appendix B Extended Dataset Analysis
 For the six representative character understanding tasks studied in our work, we selected one dataset for each category.
@@
 Love and Violence Summarization TV Shows
-Table 6: Overview of 14 tasks in the character understanding benchmark.
-Appendix D Prompts for GPT4
-for personality traits extraction ’you are a psychology research assistant designed to help analysis character personalities according to conversations.’ conversation ’Here ends the conversation.Give me a character description for each main character according to this conversation.’ Character Description Here is a character description:Monica: Responsible, caring, and organized.
+Table 6: Overview of 8 more tasks in the character understanding benchmark.
+Appendix C Prompts for GPT4
+The following is the prompt we sent to GPT4: For personality traits extraction ’you are a psychology research assistant designed to help analysis character personalities according to conversations.’ conversation ’Here ends the conversation.Give me a character description for each main character according to this conversation.’
+Character Description Here is a character description:Monica: Responsible, caring, and organized.
 Assertive and confident in her actions.Chandler: Witty and self-deprecating with an approachable sense of humor.
@@
 His behavior is indicative of someone who is going through emotional turmoil.’
-Appendix E Name Replacement and Character Comparison
-E.1 Cross-Language Uncommon Name Replacement
+Appendix D Name Replacement and Character Comparison
+We also provide partial name replacement dictionary in table  7.
+D.1 Cross-Language Uncommon Name Replacement
 With gender swap:[Bojing, Cuixia, Jingjing, Yunsheng, Meilin, Yusong] Without gender swap:[Cuixia, Bojing,Yunsheng,Jingjing,Yusong, Meilin]
@@
 Fleeing her wedding because of a sudden realization shows she can make drastic decisions based on her emotions.
-E.2 Same-Language Replacement
+D.2 Same-Language Replacement
 With Gender-Matched replacement, we replace all the character name [Monica, Joey, Chandler, Phoebe, Ross, Rachel]
@@
 With Gender Swap as well, we got [Andrew, Sally,Olivia, Ethan, Lauren,Benjamin].
-E.3 Anonymized Speaker Replacement
+D.3 Anonymized Speaker Replacement
 replace all the character name [Monica, Joey, Chandler, Phoebe, Ross, Rachel]
@@
 We use the prompt of "you are a psychology research assistant designed to help analysis character personalities according to conversations……..Here ends the conversation.Give me a character description for each main character according to this conversation." " and we get such result from TVshow Friends scene1.
-E.4 Control Test
-Example Here is a dialogue between 2 people: Sheldon : (now lifts her head) phoebe!!
-Amy : you know it’s me?
-Sheldon : for like a half an hour!
-man, you can lie about sweden!
-Amy : how can you come here?
-Sheldon : how could you not tell me you worked here?
-Amy : i don’t have to tell you everything!P2: We got an eight-point-two.
-Trust me, you’re happy.
-Prompt The following is a dialogue between the same two people. Please identify who they are:
-P0 : (singing) ïpan stripan, glupi glabi!
-änd that is the swedish national anthem!
-thank you for asking!
-(looks annoyed) P1 : wow, ikea… what a rich culture.
-uhm, you know what?
-i have a friend who is a masseuse.
-P0 : oh, ja!
-ja!
-P1 : yah!
-she’s… uhm… not very good though… (phoebe looks devastated)
-Table 8: An example for the control test, where in the example,the script is from Friends but speakers are annoted as Amy and Sheldon(from The Big Bang Theory).
-Appendix F Verbatim avoid Reasoning
-Figure 6: In a personality understanding task where the scenario involves ’Joey returned the lost wallet to its rightful owner, Monica,’ the question posed might be, ’According to this action, his personality should be: ?’ The human reasoning process would draw on contextual evidence, concluding that returning lost property demonstrates honesty.
-In contrast, the LLM might conclude, ’Joey and Monica are from the TV show Friends, and Joey is characterized as an honest character in this show,’ thus relying heavily on memorized information rather than genuine reasoning.
-Appendix G Benchmark Datasets
+Appendix E Benchmark Datasets
 The six selected benchmarks cover a range of character-centric tasks, designed to test different aspects of character understanding(Table 1):
   • TVShow Character Guessing: Based on the TVSHOWGUESS dataset, which contains scripts from five popular TV series: Friends, The Big Bang Theory, The Office, Frasier, and Gilmore Girls Sang et al.
-(2022c), this is a multi-choice classification task.
+(2022b), this is a multi-choice classification task.
 The dialogue is anonymized using placeholder identifiers (e.g., P1, P2), and the model must predict the correct speaker for each line, testing its ability to identify characters based on context.
@@
 The model is required to generate a coherent summary from dialogue utterances and scene descriptions, capturing key plot points while excluding irrelevant details.
-Appendix H Gist Prompting Strategies Across Tasks
+Appendix F Gist Prompting Strategies Across Tasks
 We design Gist Prompting methods to reduce reliance on memorization and promote inference-based character understanding. Below, we detail the task-specific prompt formulations.
-H.1 Task-Specific Gist Prompts
+F.1 Task-Specific Gist Prompts
 Character Guessing